### v1

# Mamba-based Deep Learning Approaches for Sleep Staging on a Wireless Multimodal Wearable System without Electroencephalography

Abstract
Study Objectives
We investigate using Mamba-based deep learning approaches for sleep staging on signals from ANNE One (Sibel Health, Evanston, IL), a minimally intrusive dual-sensor wireless wearable system measuring chest electrocardiography (ECG), triaxial accelerometry, and temperature, as well as finger photoplethysmography (PPG) and temperature.
Methods
We obtained wearable sensor recordings from 360 adults undergoing concurrent clinical polysomnography (PSG) at a tertiary care sleep lab. PSG recordings were scored according to AASM criteria. PSG and wearable sensor data were automatically aligned using their ECG channels with manual confirmation by visual inspection. We trained Mamba-based models with both convolutional-recurrent neural network (CRNN) and the recurrent neural network (RNN) architectures on these recordings. Ensembling of model variants with similar architectures was performed.
Results
Our best approach, after ensembling, attains a 3-class (wake, NREM, REM) balanced accuracy of 83.50%, F1 score of 84.16%, Cohen’s of 72.68%, and a MCC score of 72.84%; a 4-class (wake, N1/N2, N3, REM) balanced accuracy of 74.64%, F1 score of 74.56%, Cohen’s of 61.63%, and MCC score of 62.04%; a 5-class (wake, N1, N2, N3, REM) balanced accuracy of 64.30%, F1 score of 66.97%, Cohen’s of 53.23%, MCC score of 54.38%.
Conclusions
Deep learning models can infer major sleep stages from a wearable system without electroencephalography (EEG) and can be successfully applied to data from adults attending a tertiary care sleep clinic.
1University of Toronto Dept. of aMedicine, bMedical Biophysics, cComputer Science Toronto, ON, Canada
2Sunnybrook Research Institute aDept. of Medicine Neurology Div., bPhysical Sciences Toronto, ON, Canada
3Vector Institute, 4University Health Network Toronto, ON, Canada
5Queen’s University Dept. of Electrical & Computer Engineering Kingston, ON, Canada
[…]
andrew.lim@utoronto.ca
Keywords: Deep Learning Sleep Staging Wearable Devices
1 Introduction
1.1 Sleep Staging and Polysomnography
Sleep is essential to physical and mental health. Quantitative evaluation of sleep architecture and physiology is central to the diagnosis of many sleep disorders and plays an important role in the diagnosis of neurodegenerative diseases [1]. Sleep is traditionally and most reliably measured by polysomnography (PSG), in which brain activity, eye movements, heart rate (HR), muscle tone, and breathing patterns are recorded simultaneously using sensors attached to a patient sleeping in a laboratory environment. This is followed by sleep staging, the manual categorization of time periods into different stages of sleep based on a trained sleep technician’s visual inspection of the collected physiological data. The exact sleep staging methodology depends on the specific framework; under the American Academy of Sleep Medicine’s (AASM) [2] conventions, recordings are divided into 30-second windows, or "epochs", with electroencephalography (EEG), electromyography (EMG), and electro-oculography (EOG) used to classify epochs into three broad stages: awake, rapid eye movement (REM) sleep, and non-REM (NREM) sleep, which can be further subcategorized as N1, N2, and N3 sleep. EEG is the principal signal used for sleep staging [2, 3, 4].
1.2 Wearable Devices and Automated Sleep Staging
The need for specialized personnel, equipment, and laboratory space makes PSGs expensive to conduct and difficult to scale. The cumbersome sensor setup and unfamiliar resting environment may also be uncomfortable for many patients and can result in sleep that is unrepresentative of that of the home environment. Furthermore, the throughput of PSGs is limited by the time-consuming process of monitoring patients throughout the night and then visually labelling sleep epochs.
In response to these challenges, wearable devices have been developed that record some but not all of the signals recorded by conventional PSG. Devices recording EEG have the advantage of directly measuring brain activity but the disadvantage of the discomfort associated with sensors on the head, similar to PSG. On the other hand, EEG-less devices are less obtrusive, and more easily set up for in-home use without a technician, but infer rather than directly measure sleep stage.
There has been considerable interest in pairing wearable devices with automated sleep stage inference in order to minimize the need for specialized technologists for labeling and hence improve scalability. Indeed, even for the analysis of conventional in-labratory PSG, automated sleep staging has been explored as a means of addressing the time-intensiveness of sleep epoch labelling. The focus on pattern recognition and loosely rule-based nature of sleep staging is ideal for machine learning approaches which have employed model architectures like convolutional neural networks (CNNs) [5], recurrent neural networks (RNNs) [6], hybrids of the previous two (e.g. convolutional recurrent neural networks (CRNN)) [7], Transformers [8], and graph neural networks [9]. Deep learning models trained on PSG data, most notably the EEG signal or some combination of EEG, EMG, and EOG, have surpassed inter-scorer agreement levels, and can make predictions orders of magnitude faster than a human evaluator at the cost of sacrificed interpretability [10, 11].
1.3 Recurrent Neural Networks and Mamba
RNNs are a class of neural networks that take in a sequence of vectors and produce a vector output per input, allowing them to execute a wide breadth of time-series tasks. Prevalent modern RNN architectures include long short term memory (LSTM) [12] and gated recurrent units (GRU) [13], which are known to forget very long-range dependencies and whose architecture cannot take advantage of GPU acceleration, limiting the training speed.
Mamba [14] provides a solution to both problems as a type of state space model (SSM) [15] that combines the excellent long-range dependency modeling of SSMs with an additional time-dependent selection mechanism that allows for focusing on or ignoring particular features, as well as a hardware-friendly algorithm that provides computational efficiency. Empirically, Mamba has achieved state-of-the-art performance on supervised and unsupervised metrics using datasets of text, audio waveforms, and DNA sequences [14], while Mamba-based models have been adopted to perform automatic sleep staging using EEG and other PSG signals to high degrees of accuracy [16, 17].
![[Uncaptioned image]](extracted/6084238/figures/anne.png)
| Statistic | Range / Count | ||
| Recordings | 357 | ||
| Male/Female | 181/176 (51/49%) | ||
| Awake [%] | 43.43 | 17.15 | 11.54 95.24 |
| N1 Sleep [%] | 10.67 | 6.13 | 1.05 33.13 |
| N2 Sleep [%] | 27.18 | 10.73 | 0.58 63.59 |
| N3 Sleep [%] | 10.38 | 7.94 | 0.00 54.83 |
| REM Sleep [%] | 8.34 | 5.81 | 0.00 28.80 |
| Age [y] | 57.49 | 17.56 | 18.17 92.07 |
| Apnea-Hyponea Index (AHI) | 11.54 | 17.21 | 0.00 127.00 |
| Total Sleep Time (TST) [h] | 5.10 | 1.46 | 0.66 9.16 |
| Body Mass Index (BMI) [kg ] | 27.81 | 5.73 | 10.00 50.40 |
| Periodic Limb Movement Index (PLMI) | 15.01 | 26.99 | 0.00 169.50 |
| Sleep Efficiency [%] | 71.83 | 17.91 | 11.60 98.50 |
1.4 Contribution
This paper applies Mamba-based deep learning approaches to the problem of sleep staging using data from the ANNE One [18] sensor. ANNE One, as seen in Figure 1A, is a flexible, minimally-intrusive, and clinical-grade wireless dual-sensor wearable system by Sibel Health composed of a finger module, which measures photoplethysmography (PPG) and temperature (Temp.), and a separate module attached to the chest via an adhesive, which measures electrocardiography (ECG), triaxial accelerometry (Acc.), and temperature. The raw signals of a typical recording are plotted in Figure 1B.
An anterior study centered on ANNE One [19] exists, which primarily emphasized the development of person-specific models which were trained and evaluated on data from the same patient, with a population model trained and evaluated on data from different patients showing much poorer performance. We aim to develop a better population model that allows accurate sleep staging from ANNE One signals recorded from previously unseen patients.
There is a considerable body of work applying machine learning to the problem of sleep stage inference from non-EEG signals. For the most part, these have been trained and tested using the cardiopulmonary sensors of conventional PSG [20, 21, 22, 23, 24, 25, 26], which are typically artifact-free. Relatively few studies have applied these approaches to data from wearable devices, which are in general more prone to poor signal quality and signal loss; consequently, such studies often need to reject epochs of poorer-quality signal prior to model training. Moreover, these works have focused on analysis of data from single sensors, typically ECG or PPG [27, 28]. Finally, models developed for wearable sensors have largely trained on data from relatively young and healthy volunteers, rather than patients presenting to sleep clinics, raising a question of generalizability to clinical populations.
The current study builds on and extends this body of work in several ways. First, we leverage the ANNE One devices’ multi-sensor nature, which may in principle allow for predictions even in settings where one or more sensors are reading low-quality signals. Second, we train and test on data from relatively older patients with a range of sleep disorders (Figure 1C) rather than younger and healthier patients, ensuring applicability to older adults with sleep complaints. Finally, this work broadens the Mamba architecture to EEG-less sleep staging, and aims to answer whether the incorporation of Mamba offers a relative performance increase similar to PSG EEG sleep staging as noticed in [16, 17].
2 Method
2.1 Dataset
ANNE wearable sensors were used to collect full-night recordings from 360 adults undergoing concurrent clinical PSG at a tertiary care clinical sleep laboratory in Sunnybrook Health Sciences Centre, Toronto, Canada. PSG recordings were recorded according to AASM practice parameters for in-labratory PSGs using a Grael PSG system (Compumedics, Victoria, Australia). The study was conducted in accordance with the latest version of the Declaration of Helsinki. All subjects provided written informed consent for the full study protocol.
Participants were first instrumented with the ANNE One, whose recordings were started ahead of the set-up and initiation of PSG recordings. In order to mimic what would occur in a home environment, the ANNE signals were not monitored in real time, and no attempt was made to correct sensor-related issues (e.g. poor skin contact) overnight. Upon completion, each PSG recording was manually labelled by one of four technologists according to AASM standards [2]. In addition, the portion of the ANNE recordings preceding the onset of PSG recording was labeled as wake based on direct observation of participants by study staff. PSG and wearable sensor data were automatically aligned by their shared ECG signals using cross-correlation analysis and verified by visual inspection.
Three participants who did not sleep at all were excluded from analysis, leaving a total of 357 patients. The characteristics of the study’s participants are outlined in Figure 1C.
The machine learning task is thus defined from the dataset as multiclass 1D signal segmentation, wherein the ANNE One signals constitute the features, the PSG-annotated sleep stages form the labels, and class predictions are made per point in time, as demonstrated in Figure 2A.
2.2 Preprocessing
The ECG data were recorded at 512 Hz and preprocessed as follows. A wavelet transform was applied to raw ANNE ECG signals to identify temporal windows containing QRS complexes; within each window, raw ECG signals were searched for local maxima and minima to identify the precise timing of the R-peak. The wavelet-transformed ECG data were subsequently downsampled to 100 Hz for additional analyses. Meanwhile, the Hilbert transform was applied to extract the envelope of each QRS complex, the area under which being taken as a measure of the signal to noise ratio (i.e. signal quality index (SQI)). For each peak, morphological features (e.g. amplitude, sharpness) were extracted and hierarchical density-based spatial clustering of applications with noise (HDBSCAN) [29] served to group all identified peaks into a small number of clusters. Peaks occurring in trains with relatively stable RR intervals in the physiological range of human HR were identified as likely representing physiological QRS complexes. QRS complexes that clustered morphologically with these peaks, irrespective of regularity, were used to compute beat-to-beat heart rate.
The PPG data were recorded at 128 Hz and preprocessed as follows. First, PPG signals were band-pass filtered between and downsampled to 100 Hz. Trains of physiological pulse waves were identified on the basis of relatively stable HR and amplitude. These high-quality pulse waves were normalized to a standard amplitude and duration, and clustered using HDBSCAN. For each cluster, the mean of its normalized pulse waves was taken as a template wave for that cluster; beat-to-beat template matching was performed to determine the extent to which each pulse wave in the rest of the recording matched one of the template waves, with the lowest sum-squared similarity to the template constituting a measure of signal quality for each PPG pulse wave. For each matching pulse wave, the timing of the foot and peak was noted, from which a measure of heart rate was obtained.
The pulse arrival time (PAT) is the delay from the R-peak of the ECG to the arrival of the pulse wave in the PPG. To compute PAT from the ANNE data, the time-series of valid ECG peaks and PPG feet were matched to each other by empirically considering all possible lags within the range msec and selecting that which successfully matched the greatest number of ECG peaks and PPG feet. Afterwards, for each ECG R-peak and its corresponding PPG foot, the PAT of that heartbeat was computed, generating a beat-to-beat time series of PAT.
The chest XYZ accelerometry data were recorded at 210 Hz and low-pass filtered, followed by the computation of the euclidean norm minus one (ENMO), theta (body pitch), phi (body roll) and the z-angle (the deviation from vertical of a vector perpendicular to the face of the sensor).
2.3 Model
Two Mamba-based deep neural network architectures were investigated: a convolutional-recurrent neural network (CRNN) and a recurrent neural network (RNN) as summarized in Figures 2AB, 3. The CRNN is a typical sleep stage architecture [21, 11] that emulates the way human evaluators label PSGs: the convolutional component encodes physiological signals within each epoch while the recurrent component captures transitional relationships between sleep stages across epochs. The RNN architecture is a lightweight and parameter-scarce alternative to the CRNN, consisting of less parameters (580K vs. 70K). Since Mamba is designed for pattern recognition across long contexts, it could potentially interact with signals more effectively than the CRNN if the signals are given directly on a finer time-resolution without windowing the data.
The CRNN architecture has three input channels for time-series, frequency, and scalar data (see Figure 3): the first two are 3D inputs, with axes for the number of 30-second windows, the window size (i.e. 30 seconds worth of sequential data), and number of features, while the 2D scalar inputs have axes for the number of windows and number of features. The model comprises:
1)
A convolutional component, consisting of 1D residual blocks that sequentially downsize the time-series and frequency datas’ window size axes and flatten them into an embedding vector per window. The time-series embedding, frequency embedding and scalar data are then concatenated across the feature axis to form an embedding vector per window.
2)
A recurrent component, consisting of a bidirectional Mamba block that processes the embedding vectors sequentially, followed by an element-wise multi-layer perceptron (MLP) that diminishes the feature-size axis into logits for each class. The bidirectional Mamba block (Appendix A) is the concatenation of two Mamba blocks that process epochs in opposite directions.
The RNN architecture (Figure 2B) processes a single 2D input channel with axes for the number of -second windows and number of features, where is an arbitrary value dividing 30. It comprises:
1)
An element-wise MLP with hidden layer dimensions that downsizes the data to dimensions, where the choices of and are arbitrary.
2)
A bidirectional Mamba block that processes windows sequentially.
3)
An element-wise MLP that downsizes the data into logits, which, during inference-time, can be grouped and averaged to form logits per 30-second window to comply with AASM standards.
Our previous investigations [30] (Appendix C) have found engineered frequency domain and scalar features to yield significant performance benefits for GRU-based deep learning approaches. The engineered scalar features distribute differently depending on sleep stage, while changes to their frequency domains often designate sleep stage transitions (Appendix B). Thus, empirical selections of these features (Table 1) were introduced alongside the time-series signals for both the CRNN and RNN.
A series of transformations were applied to further adjust the preprocessed data to the individual architectures. For the CRNN, all signals were resampled to 25 Hz and grouped to 30-second windows, resulting in data points per window. Then, depending on feature type (see Table 1),
For time-series features, the signal was left alone.
For frequency features, the log-magnitude spectrum was taken of the detrended signal over the 30-second axis.
For scalar features, the signal was reduced over the 30-second axis. Metadata, in the form of patient age and sex (male = 0, female = 1), was incorporated into the scalar features by repeating the constants times throughout the recording.
For the RNN, the accelerometry, ECG, and PPG signals were detrended, with the accelerometry signals further band-pass filtered over Hz. Log-magnitude spectrograms were taken with a window size of 100 points and no overlap, then finally concatenated with limb/chest temperature, which was resampled to 1 Hz, and sleep stage label, of which each point was repeated 30 times to 1 Hz.
RNN CRNN Time-Series Frequency Scalar ECG* Heart Rate (ECG + PPG) ECG* Age PPG* Heart Rate (ECG) PPG* Sex X* Heart Rate (PPG) X* Mean, ECG SQI Y* Pulse Arrival Time (PAT) Y* Mean, PPG SQI Z* Euclidean Norm Minus One (ENMO) Z* Entropy, PPG Chest Temp. Z-Angle Z-Angle* Entropy, PPG* Limb Temp. Limb-Chest Temperature Difference Skew, PPG* Theta Scale, Heart Rate (ECG + PPG) Phi Kurtosis, ENMO Entropy, Z-Angle
2.4 Training
The training, validation, and testing sets were randomly split in a size ratio to sizes . During the training process, recordings were clipped to random segments to prevent the model from overfitting to the high-level structure of most recordings, which mostly consisted of wake periods at the start and end of the recording with sleep in the middle. Cross-entropy loss was weighted on inverse class frequency to adjust for the moderate class imbalance as encapsulated by Figure 1C. The model from the iteration that maximized balanced validation accuracy was chosen as the final model. Both CRNN and RNN models were trained with the Adam optimizer [32].
The CRNN model was trained with single batches. Recordings were clipped to segments of size varying uniformly between and , and age/sex values individually were set to with probability , an arbitrary value, as a form of data augmentation. The RNN model was trained on a batch size of , with recordings clipped randomly to sequence lengths of seconds.
Three models of varying sleep stage resolution were built: a 3-class, 4-class, and 5-class sleep staging model. The 5-class model has classes for wake, N1, N2, N3, and REM sleep, while the 3-class model groups N1, N2, and N3 sleep into non-REM sleep. The 4-class model designates N1 and N2 sleep as "light sleep" as opposed to N3 or "deep sleep", a common practice in automatic sleep staging [4].
2.5 Ensembling
For each of the classes inside a -class ensemble model, a binary model was trained to predict the presence or absence of that class. The logits of these binary models were stacked with logits from the regular -class models to form the inputs to a suite of decision tree-based architectures including AdaBoost [33], random forests [34], extra-tree classifiers [35], histogram gradient boosting [36], and XGBoost [37]. The model among the suite that maximized balanced validation accuracy was selected as the final ensembler model. A visual representation of the ensembling pipeline is presented in Figure 2C.
As decision tree-based architectures do not output class-wise logits, epoch-wise predictions on the RNN model were generated by aggregating 30 one-second predictions and reducing with the mode.
3 Results
Results are compiled of 3-class, 4-class, and 5-class sleep staging for the RNN and CRNN architectures and for regular and ensemble models. Overall performance is evaluated by aggregating all test set recordings and computing the balanced accuracy (B.Acc.), precision (Prec.), recall (Rec.), F1, Cohen’s Kappa (), and Matthews correlation coefficient (MCC). We additionally conduct analyses of averaged per-recording metrics, biases in the amount of each sleep stage, and differences in performance across age, PLMI, AHI, BMI, TST, and sleep efficiency. Because individual recordings do not necessarily contain instances of all five sleep stages, balanced accuracy may not be a sensible metric for recording-wise comparisons; hence, accuracy (Acc.) is presented in lieu of balanced accuracy for those cases.
3.1 Model Evaluation
Typical-case recording predictions for all 5-class models are shown in Figure 4 along with predicted point-wise output class probabilities for models that can produce them. A Uniform Manifold Approximation and Projection (UMAP) [38] of an intermediate CRNN output is also provided in Appendix E, which shows the separation of sleep stages as clusters in the learned representation space.
Figure 5 displays the confusion matrices for all -class ensemble models and overall metrics across all -class regular and ensemble models. No significant deviations in performance were observed in the RNN models as tested on 1-second windows and as tested on 30-second windows, of which the latter is presented in this paper. The RNN ensemble model yields the best metrics and dominates most class-wise metrics (Appendix D). Ensemble models perform better than their single-model counterparts, while RNNs perform better than the CRNNs despite the shallower feature set and smaller parameter count. Most misclassifications are made between N1 and every other class, as well as between N2 and N3 for 5-class sleep staging.
Lastly, in a set of young and healthy sleepers (age < 40, AHI < 5, PLMI < 5), the RNN ensemble model’s performance is significantly boosted as denoted in the confusion matrices and metrics of Figure 6.
![[Uncaptioned image]](extracted/6084238/figures/ensemble_outputs.png)
[…]
![[Uncaptioned image]](extracted/6084238/figures/healthy-subset.png)
Metric (%) Classes Balanced Accuracy Precision Recall F1 Cohen’s MCC 5 69.26 76.85 70.07 71.55 60.48 61.60 4 79.39 78.82 77.00 76.89 67.23 67.92 3 86.32 87.63 86.70 86.88 77.87 78.14
3.2 Global Sleep Metrics
As a means of evaluating model bias and verifying model robustness, Figure 7 lays out scatter plots and Bland-Altman plots for the 5-class RNN ensemble, comparing RNN predicted vs. PSG actual values for a number of summary metrics including time asleep, % N1 sleep,% N2 sleep, % N3 sleep, and % REM sleep. The model trends towards slightly underestimating N2 sleep and slightly overestimating N1 sleep and wake.
3.3 Performance by Clinical Group
Averaged recording-wise metrics are cataloged in Table 2 for the 5-class ensemble CRNN and RNN, wherein linear regressions were performed to fit various recording metrics to accuracy. Performance of the RNN ensemble is robust to differences in sex, periodic limb movement index, body mass index, total sleep time and sleep efficiency, but shows slightly worse performance with greater age and higher AHI. The CRNN model is robust to differences in age, sex, PLMI, and BMI but shows slightly worse performance with higher AHI, shorter TST, and lower sleep efficiency. Notably, even where performance varies by clinical factors, the percentage of explained change in performance is nevertheless weak for all metrics (). There is no significant difference in accuracy between the raters in the testing set (RNN , CRNN , one-way F-test).
| RNN | CRNN | ||||||
| Metric (%) | |||||||
| Accuracy | 65.09 | 9.84 | 62.58 | 12.50 | |||
| F1 | 67.69 | 9.92 | 63.81 | 14.17 | |||
| Cohen’s | 50.39 | 14.91 | 46.33 | 16.13 | |||
| Matthews Correlation Coefficient | 52.61 | 13.94 | 48.77 | 15.91 | |||
| Sleep Statistic | |||||||
| Age | 0.0827 | 0.0350 | 0.0346 | 0.1779 | |||
| Sex | 0.0689 | 0.0552 | 0.024 | 0.2630 | |||
| Periodic Limb Movement Index (PLMI) | 0.0120 | 0.4300 | 0.0012 | 0.8061 | |||
| Apnea-Hypopnea Index (AHI) | 0.1525 | 0.0350 | 0.1622 | 0.0025 | |||
| Body Mass Index (BMI) | 0.0164 | 0.3559 | 0.0000 | 0.9824 | |||
| Total Sleep Time (TST) | 0.0000 | 0.9921 | 0.0853 | 0.0322 | |||
| Sleep Efficiency | 0.0045 | 0.6316 | 0.1503 | 0.0038 | |||
4 Discussion
4.1 Key Findings
This work has successfully developed Mamba-based CRNN and RNN models to accurately perform inference from the ANNE One device for major sleep stages. The models demonstrated good performance across a wide range of patient characteristics, without a need to screen and exclude recordings and epochs on the basis of signal quality, and with the RNN architecture slightly outperforming the CRNN architecture.
4.2 Related Studies
Previous studies have undertaken the challenge of inferring sleep stage from non-EEG signals (Tables 3 & 4) using a mixture of cardiopulmonary (ECG, PPG, respiratory impedance bands, respiratory flow cannulas) and accelerometry data. Many studies used data from conventional PSG recordings [20, 21, 22, 23, 24, 25, 26] although a handful of studies used data from purpose-designed wearable sensors [28, 21, 19, 23, 24, 27]. In general, models trained on and applied to PSG data performed better than models utilizing wearable sensor data. This is not surprising since signal quality is often much higher for attended in-laboratory PSGs where more intrusive sensors can be used, and signal quality is continuously monitored and issues can be addressed by the technologist as they occur. Even so, despite using wearable rather than PSG sensor data, the performance of our ensemble RNN model approached [21, 24, 25, 26, 20, 22] or exceeded [23] that of models trained on much larger sets of PSG data.
Our model exceeded the overall performance of most wearable sensor based models [27, 23, 28] with the exception of one [24]. Of note, this study examined a younger set of participants (mean age 45) who were free of psychiatric and neurological co-morbidity. In contrast, we studied a much older (mean age 57.49) participants with significant sleep co-morbidity (45.81% with AHI > 5). When we limit our analyses to younger individuals without sleep apnea or significant periodic limb movements, model performance improves further (Figure 6). Another contrast between [24] and the present study is that [24] performed much stricter quality control for recording quality with 54/136 or nearly 40% of nights of wearable sensor recording excluded from analysis due to poor signal quality. In contrast, in the interest of evaluating real-world performance, we did not exclude any recording or any epochs on the basis of signal quality and the reported results reflect consideration of all available epochs of recording, irrespective of signal quality. The robustness of our model to missing data or poor signal quality may be related in part to its use of multiple streams of data such that sensors returning poor quality data are compensated for in the model by those returning good quality data.
Of particular comparative interest is the one previous paper addressing sleep stage inference from ANNE One data [19]. Ultimately, success in developing a population model was limited, and while there was some success in generating person-specific models, the performance of our ensemble RNN model exceeded even the person-specific models generated in this paper.
[…]
4.3 Model Performance
Both model architectures obtain maximal per-class performances for wake and REM, and minimal per-class performances for N1 and N2 sleep. As a transitory state, periods of N1 sleep are short and instable, making its discrimination a difficult task. Physiological similarities between N2 and N3 sleep, such as cardiorespiratory signatures, similarly complicate the model’s ability to differentiate them, as evidenced by N2-N3 confusion as the most frequent misprediction for 5-class models (as well as N1/N2-N3 confusion for 4-class models). Conversely, wake and REM sleep have clearer differentiating characteristics, such as higher heart rate, which allows for easier classification.
In spite of a simpler feature set and smaller parameter count, the RNNs outperform the CRNNs in virtually every metric. It is possible that the independent processing of segmented windows by the CNN component interfered with the wider CRNN’s ability to discern changes at neighbouring window boundaries. As well, during epochs where sleep is transitioning between classes, the provided label becomes less reliable; the CRNN attempts to predict a behaviourally inconsistent epoch while it is likely the RNN more flexibly captures this ambiguity by means of multiple differing fine-resolution class predictions throughout the 30 seconds.
No statistically significant variations in performance based on PLMI, BMI, TST, or sleep efficiency were exhibited by the RNN 5-class ensemble model, providing an important advantage when applied in clinical contexts. The significant () association discovered between AHI and accuracy may owe to the fact that individuals with sleep apnea experience sleep fragmentation, resulting in more time in N1 sleep, one of the most-difficult stages to separate; even so, the association was modest (). Model performance also correlates between age and accuracy (), though this may in part be explained by the weak agreement between AHI and age ().
4.4 Limitations, Strengths, and Future Directions
A few methodological limitations are worth considering. 1) The differing nature of the way our CRNN and RNN models handle preprocessing and predictions (notably, the 30-second window employed by the CRNN as opposed to the 1-second window used in the RNN) make the models difficult to compare; our reconciling of the RNN to make predictions for 30-second windows may not reflect its full performance. 2) As neural networks, CRNNs and RNNs provide limited insight into the decoding of sleep stages from the given sensor data. Further analyses targeted toward interpreting the neural networks may be explored. 3) The recordings were overnight with a brief period of wakefulness before lights out, resulting in a preponderance of sleep vs. wake. It is difficult to know for sure how the model will perform in detecting daytime naps during long ambulatory recordings and further studies with ambulatory recordings in clinical populations will be informative.
This study also had many strengths: 1) The ANNE One system itself, with multiple channels of data recorded from a pair or minimally intrusive sensors, is a strength insofar as it provides an element of redundancy should one sensor return low quality signal, 2) the combination of Mamba with an RNN architecture allowed the achievement of relatively high performance with a relatively parameter-scarce architecture, and 3) the study population was varied, including participants from across a spectrum of sleep disorders.
5 Conclusion
This work introduces CRNN and RNN models to automatically perform 3-class, 4-class, and 5-class sleep staging using data from the ANNE One dual sensor using sleep stages inferred from concurrent PSG recordings, with the RNN ensemble model topping the others in virtually every performance metric. The model meets or surpasses metrics in related literature, especially for wearable devices, and demonstrates robustness across a wide range of clinical characteristics. Moreover, the model’s capability was achieved without prescreening and excluding large numbers of patients or epochs for reasons of signal quality. Compared to PSGs, the ANNE One sensor is far less intrusive of a device and procedure, which is significant for expanding the accessibility of sleep measurement for older and remote-living adults at greater risk of sleep-related illness. Our model provides a scalable, accurate minimally-intrusive approach to the ambulatory assessment of sleep staging without the need for EEG.
Acknowledgments
This research is supported by the Centre for Aging and Brain Health Innovation, Canadian Institutes of Health Research, and National Institute on Aging.
This research is not funded by Sibel Health, the manufacturer of the ANNE One device.
Special thanks to Prof. Jimmy Ba from the Department of Computer Science; Prof. Paul Kushner and Prof. Dylan Jones from the Department of Physics at the University of Toronto for offering various forms of support for this research during its infancy.
Appendix A: Bidirectional Mamba Block
Appendix B: Scalar and Frequency Features
Appendix C: Explorations in Feature Selection
A previous iteration of the model used a pure CRNN (GRU) approach, and the approach encompassed two additional strategies, which are utilized within the paper without extensive justification.
The engineered features are a suite of additional features across the time, frequency, and scalar domains used alongside the raw signals as extracted from the ANNE sensors.
Time Domain Frequency Domain Scalar Domain PPG, ECG Heart Rates PPG Freq. Spectra PPG Freq. Spectra Skew PPG, ECG Signal Quality Index ECG Freq. Spectra PPG Time Entropy PPG, ECG Amplitude x-accel. Freq. Spectra PPG Freq. Spectra Entropy Pulse Arrival Time (PAT) y-accel. Freq. Spectra | Heart Rate| Accel. ENMO z-accel. Freq. Spectra Accel. ENMO Kurtosis SpO2 Accel. z-angle Freq. Spectra Entropy Metadata (Gender, Age)
In addition, the auxiliary training objective was introduced in the form of the transition state. These states are derived directly from the target labels. With this, the model attempts to predict whether there will be a change in sleep class from this epoch to the next. The rationale is that sleep stages in reality do not always change at the end of 30-second windows, and many epochs may have data that is transitioning between two separate sleep stages, which would be difficult for the model as they may not resemble other members of their labelled class and instead as a mixture of both sleep stages. By training the model to also predict correct transitions, it should assist the model in determining which epochs may be occurring during a transition between stages, and thus make it more robust in general.
The application of the transition state as an auxiliary objective was attempted in the Mamba approaches in the main paper; however, it did not yield a noticeable increase in performance, and thus was not included in the final models.
The results of the approach are detailed in the graph below, with results indicating that the two approaches mentioned positively impacted results.
3-class Sleep Staging 2-class Sleep Detection F1 Precision Accuracy F1 Precision Accuracy Wake Time Series Only Baseline 0.681 0.723 0.644 0.740 0.649 0.860 Engineered Features 0.743 0.723 0.764 0.742 0.703 0.786 Engineered Features & Auxiliary Objective 0.727 0.757 0.700 0.781 0.701 0.880 Ensemble 0.732 0.789 0.683 NREM Sleep Time Series Only Baseline 0.534 0.720 0.425 0.793 0.892 0.714 Engineered Features 0.507 0.754 0.381 0.826 0.859 0.826 Engineered Features & Auxiliary Objective 0.619 0.744 0.530 0.834 0.912 0.769 Ensemble 0.702 0.767 0.647 REM Time Series Only Baseline 0.401 0.265 0.827 Engineered Features 0.428 0.287 0.844 Engineered Features & Auxiliary Objective 0.447 0.314 0.778 Ensemble 0.517 0.381 0.679 Macro Time Series Only Baseline 0.539 0.632 0.555 0.767 0.771 0.770 Engineered Features 0.559 0.663 0.581 0.783 0.784 0.792 Engineered Features & Auxiliary Objective 0.600 0.598 0.624 0.807 0.807 0.811 Ensemble 0.651 0.646 0.679
Appendix D: Class-Wise Evaluation Metrics
5-Class Sleep Staging 4-Class Sleep Staging 3-Class Sleep Staging B.Acc. Prec. Rec. F1 MCC B.Acc. Prec. Rec. F1 MCC B.Acc. Prec. Rec. F1 MCC Wake Wake Wake CRNN 83.52 84.10 84.13 84.07 67.46 67.53 84.59 84.80 84.75 84.77 69.01 69.02 84.57 84.70 84.45 84.50 68.57 68.70 RNN 86.00 88.85 87.59 87.29 74.01 75.63 88.00 89.57 89.08 88.94 77.36 78.05 88.67 89.26 89.23 89.18 77.90 78.01 CRNN Ens. 85.63 86.90 86.64 86.49 72.36 72.83 86.96 87.40 87.41 87.38 74.24 74.28 88.05 88.52 88.52 88.49 76.49 76.55 RNN Ens. 86.68 88.63 87.95 87.75 74.93 75.88 88.49 89.14 89.10 89.04 77.61 77.74 89.03 89.41 89.42 89.39 78.36 78.39 N1 Sleep N1/N2 Sleep Non-REM Sleep CRNN 66.02 86.62 79.98 82.65 22.47 24.23 62.09 67.98 68.51 65.69 26.58 29.20 79.70 80.20 79.98 79.88 59.68 60.00 RNN 67.77 87.06 77.97 81.43 22.61 25.40 68.42 71.65 72.13 71.34 38.39 39.00 81.68 82.36 82.00 81.89 63.72 64.18 CRNN Ens. 68.82 87.42 80.47 83.18 25.87 28.12 71.45 74.71 75.07 74.32 44.79 45.56 82.29 82.67 82.52 82.46 64.83 65.05 RNN Ens. 68.96 87.42 79.35 82.44 24.96 27.61 73.52 76.06 76.40 75.96 48.45 48.83 85.15 85.35 85.30 85.28 70.47 70.56 N2 Sleep CRNN 65.60 74.43 76.10 74.58 34.46 35.48 RNN 64.82 76.06 77.46 74.93 34.63 37.49 CRNN Ens. 65.29 75.54 77.14 75.00 35.03 37.10 RNN Ens. 68.12 77.33 78.59 77.02 40.56 42.13 N3 Sleep N3 Sleep CRNN 73.95 90.34 89.41 89.83 44.03 44.26 75.29 89.62 81.09 84.14 31.74 35.77 RNN 82.25 91.85 87.58 89.08 47.45 50.21 82.22 91.58 84.59 86.94 42.11 46.62 CRNN Ens. 77.66 91.19 89.59 90.25 48.04 48.69 78.35 91.05 88.33 89.39 45.80 47.15 RNN Ens. 81.20 92.14 90.30 91.01 52.80 53.78 82.18 92.12 89.32 90.35 51.18 52.91 REM Sleep REM Sleep REM Sleep CRNN 72.17 91.05 91.33 91.18 45.86 45.90 73.16 90.10 87.77 88.75 38.54 39.38 78.98 92.24 91.29 91.70 52.43 52.79 RNN 82.42 93.23 92.29 92.67 58.19 58.66 82.85 93.17 91.90 92.40 57.33 58.04 87.46 93.49 88.87 90.37 52.70 56.55 CRNN Ens. 81.70 92.69 91.06 91.70 54.01 54.92 83.06 92.82 90.54 91.38 53.64 55.13 83.47 92.73 89.69 90.78 51.80 53.90 RNN Ens. 84.79 94.20 93.63 93.86 64.41 64.68 87.16 94.40 93.22 93.65 64.47 65.33 88.17 94.49 92.94 93.48 61.16 65.42
Appendix E: UMAP of CRNN Model Learned Embeddings

### v2

@@
 Study Objectives
-We investigate using Mamba-based deep learning approaches for sleep staging on signals from ANNE One (Sibel Health, Evanston, IL), a minimally intrusive dual-sensor wireless wearable system measuring chest electrocardiography (ECG), triaxial accelerometry, and temperature, as well as finger photoplethysmography (PPG) and temperature.
+We investigate Mamba-based deep learning approaches for sleep staging on signals from ANNE One (Sibel Health, Evanston, IL), a non-intrusive dual-module wireless wearable system measuring chest electrocardiography (ECG), accelerometry, and temperature, and finger photoplethysmography (PPG) and temperature.
 Methods
-We obtained wearable sensor recordings from 360 adults undergoing concurrent clinical polysomnography (PSG) at a tertiary care sleep lab.
-PSG recordings were scored according to AASM criteria.
+We obtained wearable sensor recordings from 360 adults undergoing concurrent polysomnography (PSG) at a tertiary care sleep lab.
+Each PSG recording was manually scored and these annotations served as ground truth labels for training and evaluation of our models.
 PSG and wearable sensor data were automatically aligned using their ECG channels with manual confirmation by visual inspection.
-We trained Mamba-based models with both convolutional-recurrent neural network (CRNN) and the recurrent neural network (RNN) architectures on these recordings.
+We trained Mamba-based models with convolutional-recurrent neural network (CRNN) and the recurrent neural network (RNN) architectures on these recordings.
 Ensembling of model variants with similar architectures was performed.
 Results
-Our best approach, after ensembling, attains a 3-class (wake, NREM, REM) balanced accuracy of 83.50%, F1 score of 84.16%, Cohen’s of 72.68%, and a MCC score of 72.84%; a 4-class (wake, N1/N2, N3, REM) balanced accuracy of 74.64%, F1 score of 74.56%, Cohen’s of 61.63%, and MCC score of 62.04%; a 5-class (wake, N1, N2, N3, REM) balanced accuracy of 64.30%, F1 score of 66.97%, Cohen’s of 53.23%, MCC score of 54.38%.
+Our best approach, after ensembling, attains a 3-class (wake, non rapid eye movement [NREM] sleep, rapid eye movement [REM] sleep) balanced accuracy of 84.02%, F1 score of 84.23%, Cohen’s of 72.89%, and a Matthews correlation coefficient (MCC) score of 73.00%; a 4-class (wake, NREM stage 1/2 [N1/N2], NREM stage 3 [N3], REM) balanced accuracy of 75.30%, F1 score of 74.10%, Cohen’s of 61.51%, and MCC score of 61.95%; a 5-class (wake, N1, N2, N3, REM) balanced accuracy of 65.11%, F1 score of 66.15%, Cohen’s of 53.23%, MCC score of 54.38%.
 Conclusions
-Deep learning models can infer major sleep stages from a wearable system without electroencephalography (EEG) and can be successfully applied to data from adults attending a tertiary care sleep clinic.
+Deep learning models can infer major sleep stages from the ANNE One, a wearable system without electroencephalography (EEG), and can be successfully applied to data from adults attending a tertiary care sleep clinic.
 1University of Toronto Dept. of aMedicine, bMedical Biophysics, cComputer Science Toronto, ON, Canada
 2Sunnybrook Research Institute aDept. of Medicine Neurology Div., bPhysical Sciences Toronto, ON, Canada
-3Vector Institute, 4University Health Network Toronto, ON, Canada
-5Queen’s University Dept. of Electrical & Computer Engineering Kingston, ON, Canada
+3Vector Institute Toronto, ON, Canada
+4University Health Network Toronto, ON, Canada
+5Beth Israel Deaconess Medical Center Boston, MA, USA
+6Rush University Medical Center Chicago, IL, USA
+7Queen’s University Dept. of Electrical & Computer Engineering Kingston, ON, Canada
 […]
@@
 Sleep is essential to physical and mental health.
-Quantitative evaluation of sleep architecture and physiology is central to the diagnosis of many sleep disorders and plays an important role in the diagnosis of neurodegenerative diseases [1].
+Quantitative evaluation of sleep architecture and physiology is central to the diagnosis of many sleep disorders and plays an important role in evaluating individuals with and at risk for neurodegenerative diseases [1].
 Sleep is traditionally and most reliably measured by polysomnography (PSG), in which brain activity, eye movements, heart rate (HR), muscle tone, and breathing patterns are recorded simultaneously using sensors attached to a patient sleeping in a laboratory environment.
-This is followed by sleep staging, the manual categorization of time periods into different stages of sleep based on a trained sleep technician’s visual inspection of the collected physiological data.
-The exact sleep staging methodology depends on the specific framework; under the American Academy of Sleep Medicine’s (AASM) [2] conventions, recordings are divided into 30-second windows, or "epochs", with electroencephalography (EEG), electromyography (EMG), and electro-oculography (EOG) used to classify epochs into three broad stages: awake, rapid eye movement (REM) sleep, and non-REM (NREM) sleep, which can be further subcategorized as N1, N2, and N3 sleep.
+This is followed by sleep staging, the manual categorization of time periods into different sleep stages based on a trained sleep technician’s visual inspection of the collected physiological data.
+The exact sleep staging methodology depends on the specific framework; under the American Academy of Sleep Medicine’s (AASM) [2] conventions, recordings are divided into 30-second windows, or "epochs", with electroencephalography (EEG), electromyography (EMG), and electro-oculography (EOG) used to classify epochs into three stages: awake, rapid eye movement (REM) sleep, and non-REM (NREM) sleep, which can be further subcategorized as N1, N2, and N3 sleep, with N1 as the lightest and N3 as the deepest NREM sleep.
 EEG is the principal signal used for sleep staging [2, 3, 4].
@@
 The need for specialized personnel, equipment, and laboratory space makes PSGs expensive to conduct and difficult to scale.
-The cumbersome sensor setup and unfamiliar resting environment may also be uncomfortable for many patients and can result in sleep that is unrepresentative of that of the home environment.
-Furthermore, the throughput of PSGs is limited by the time-consuming process of monitoring patients throughout the night and then visually labelling sleep epochs.
-In response to these challenges, wearable devices have been developed that record some but not all of the signals recorded by conventional PSG.
-Devices recording EEG have the advantage of directly measuring brain activity but the disadvantage of the discomfort associated with sensors on the head, similar to PSG.
-On the other hand, EEG-less devices are less obtrusive, and more easily set up for in-home use without a technician, but infer rather than directly measure sleep stage.
+The cumbersome sensor setup and unfamiliar resting environment may discomfort patients, resulting in sleep that is unrepresentative of that in the home environment.
+Particularly sick or frail older adults may be physically unable to attend in-laboratory PSG, in addition to those living in rural or remote areas without PSG services.
+Furthermore, monitoring patients throughout the night and visually labelling sleep epochs is a time-consuming process that limits the throughput of sleep analyses.
+In response to these challenges, wearable devices have been developed that record a subset of conventional PSG signals.
+Devices recording EEG have the advantage of directly measuring brain activity but the disadvantage of the discomfort associated with sensors on the head similar to PSGs.
+On the other hand, EEG-less devices are less obtrusive and easily set up for in-home use without a technician, but infer rather than directly measure sleep stage.
 There has been considerable interest in pairing wearable devices with automated sleep stage inference in order to minimize the need for specialized technologists for labeling and hence improve scalability.
@@
 convolutional recurrent neural networks (CRNN)) [7], transformers [8], and graph neural networks [9].
-Deep learning models trained on PSG data, most notably the EEG signal or some combination of EEG, EMG, and EOG, have surpassed inter-scorer agreement levels, and can make predictions orders of magnitude faster than a human evaluator at the cost of sacrificed interpretability [10, 11].
+Deep learning models trained on PSG data, most notably the EEG signal or some combination of EEG, EMG, and EOG, have surpassed inter-scorer agreement levels, with 5-class accuracy and Cohen’s as high as 92% and 86% respectively, and can make predictions orders of magnitude faster than a human evaluator at the cost of sacrificed interpretability [10, 11].
 1.3 Recurrent Neural Networks and Mamba
-RNNs are a class of neural networks that take in a sequence of vectors and produce a vector output per input, allowing them to execute a wide breadth of time-series tasks.
+RNNs are a class of neural networks that takes in a sequence of vectors (e.g.
+a time series of pulse waves, or breathing) and produce a vector output per input, allowing them to execute a wide breadth of time-series tasks.
 Prevalent modern RNN architectures include long short term memory (LSTM) [12] and gated recurrent units (GRU) [13], which are known to forget very long-range dependencies and whose architecture cannot take advantage of GPU acceleration, limiting the training speed.
@@
 Empirically, Mamba has achieved state-of-the-art performance on supervised and unsupervised metrics using datasets of text, audio waveforms, and DNA sequences [14], while Mamba-based models have been adopted to perform automatic sleep staging using EEG and other PSG signals to high degrees of accuracy [16, 17].
-![[Uncaptioned image]](extracted/6084238/figures/anne.png)
 Statistic | Range / Count | ||
 Recordings | 357 | ||
-| Male/Female | 181/176 (51/49%) | ||
-| Awake [%] | 43.43 | 17.15 | 11.54 95.24 |
-| N1 Sleep [%] | 10.67 | 6.13 | 1.05 33.13 |
-| N2 Sleep [%] | 27.18 | 10.73 | 0.58 63.59 |
-| N3 Sleep [%] | 10.38 | 7.94 | 0.00 54.83 |
-| REM Sleep [%] | 8.34 | 5.81 | 0.00 28.80 |
+Male/Female | 181/176 (51%/49%) | ||
+Recording Duration [h] | 8.16 | 1.39 | 1.75 12.94 |
+Awake [%] | 40.96 | 16.58 | 10.47 95.24 |
+N1 Sleep [%] | 11.13 | 6.24 | 1.05 38.42 |
+N2 Sleep [%] | 28.37 | 10.61 | 0.58 63.59 |
+N3 Sleep [%] | 10.85 | 8.12 | 0.00 54.83 |
+REM Sleep [%] | 8.70 | 6.08 | 0.00 29.14 |
 Age [y] | 57.49 | 17.56 | 18.17 92.07 |
@@
 Total Sleep Time (TST) [h] | 5.10 | 1.46 | 0.66 9.16 |
-| Body Mass Index (BMI) [kg ] | 27.81 | 5.73 | 10.00 50.40 |
+Body Mass Index (BMI) [kg ] | 27.88 | 5.59 | 16.90 50.40 |
 Periodic Limb Movement Index (PLMI) | 15.01 | 26.99 | 0.00 169.50 |
@@
 This paper applies Mamba-based deep learning approaches to the problem of sleep staging using data from the ANNE One [18] sensor.
-ANNE One, as seen in Figure 1A, is a flexible, minimally-intrusive, and clinical-grade wireless dual-sensor wearable system by Sibel Health composed of a finger module, which measures photoplethysmography (PPG) and temperature (Temp.), and a separate module attached to the chest via an adhesive, which measures electrocardiography (ECG), triaxial accelerometry (Acc.), and temperature.
-The raw signals of a typical recording are plotted in Figure 1B.
-An anterior study centered on ANNE One [19] exists, which primarily emphasized the development of person-specific models which were trained and evaluated on data from the same patient, with a population model trained and evaluated on data from different patients showing much poorer performance.
-We aim to develop a better population model that allows accurate sleep staging from ANNE One signals recorded from previously unseen patients.
+ANNE One, as seen in Figure 1A, is a flexible, minimally-intrusive, and clinical-grade wireless dual-module wearable system by Sibel Health composed of a finger module, which measures photoplethysmography (PPG) and temperature, and a separate module attached to the chest via an adhesive, which measures electrocardiography (ECG), triaxial accelerometry, and temperature.
+The raw signals of a typical recording are plotted in Figure 1B alongside its hypnogram.
+A prior study centered on ANNE One [19] exists, which primarily emphasized the development of person-specific models which were trained and evaluated on data from the same patient, with a population model trained and evaluated on data from different patients showing much poorer performance.
+We aim to develop a superior population model to achieve accurate sleep staging from ANNE One signals recorded from previously unseen patients.
 There is a considerable body of work applying machine learning to the problem of sleep stage inference from non-EEG signals.
 For the most part, these have been trained and tested using the cardiopulmonary sensors of conventional PSG [20, 21, 22, 23, 24, 25, 26], which are typically artifact-free.
-Relatively few studies have applied these approaches to data from wearable devices, which are in general more prone to poor signal quality and signal loss; consequently, such studies often need to reject epochs of poorer-quality signal prior to model training.
+Relatively few studies have applied these approaches to data from wearable devices, which are in general more prone to poor signal quality and signal loss; consequently, such studies often need to reject epochs of poorer-quality signal prior to model training and prediction.
 Moreover, these works have focused on analysis of data from single sensors, typically ECG or PPG [27, 28].
@@
 First, we leverage the ANNE One devices’ multi-sensor nature, which may in principle allow for predictions even in settings where one or more sensors are reading low-quality signals.
-Second, we train and test on data from relatively older patients with a range of sleep disorders (Figure 1C) rather than younger and healthier patients, ensuring applicability to older adults with sleep complaints.
-Finally, this work broadens the Mamba architecture to EEG-less sleep staging, and aims to answer whether the incorporation of Mamba offers a relative performance increase similar to PSG EEG sleep staging as noticed in [16, 17].
+Second, we train and test on data from relatively older patients with a range of sleep disorders (41% were over the age of 65 and 46% had sleep apnea; Figure 1C) rather than younger and healthier patients, ensuring applicability to older adults with sleep complaints.
+This is particularly important as age, sleep, and neurological conditions can all plausibly alter the relationship between sleep stage and autonomic, cardiac, and pulmonary function, meaning that models trained on younger adults may generalize poorly to older adults.
+Finally, this work broadens the Mamba architecture to EEG-less sleep staging by exploring whether Mamba’s relative performance increases [16, 17] are also noticed in this context.
 2 Method
@@
 ANNE wearable sensors were used to collect full-night recordings from 360 adults undergoing concurrent clinical PSG at a tertiary care clinical sleep laboratory in Sunnybrook Health Sciences Centre, Toronto, Canada.
-PSG recordings were recorded according to AASM practice parameters for in-labratory PSGs using a Grael PSG system (Compumedics, Victoria, Australia).
+PSGs were recorded according to AASM practice parameters for in-labratory PSGs with a Grael PSG system (Compumedics, Victoria, Australia).
 The study was conducted in accordance with the latest version of the Declaration of Helsinki.
@@
 poor skin contact) overnight.
-Upon completion, each PSG recording was manually labelled by one of four technologists according to AASM standards [2].
+Upon completion, each PSG recording was manually labeled by one of four technologists according to AASM standards [2].
 In addition, the portion of the ANNE recordings preceding the onset of PSG recording was labeled as wake based on direct observation of participants by study staff.
 PSG and wearable sensor data were automatically aligned by their shared ECG signals using cross-correlation analysis and verified by visual inspection.
-Three participants who did not sleep at all were excluded from analysis, leaving a total of 357 patients. The characteristics of the study’s participants are outlined in Figure 1C.
-The machine learning task is thus defined from the dataset as multiclass 1D signal segmentation, wherein the ANNE One signals constitute the features, the PSG-annotated sleep stages form the labels, and class predictions are made per point in time, as demonstrated in Figure 2A.
+Three participants who did not sleep at all were excluded from analysis, as were time periods prior to the sensor placement or after the sensors were taken off, leaving a total of 357 patients with 2914 hours of recording.
+The characteristics of the study’s participants are outlined in Figure 1C.
+Generally, older participants (age 65) and those with sleep apnea had poorer sleep quality, with lower total sleep time, poorer sleep efficiency, and increased N1 sleep time relative to other sleep stages (Appendix A).
+The machine learning task is defined from the dataset as a form of multiclass 1D signal segmentation, wherein the ANNE One signals constitute the features, the PSG-annotated sleep stages form the labels, and class predictions are made per point in time, as demonstrated in Figure 2A.
 2.2 Preprocessing
@@
 The wavelet-transformed ECG data were subsequently downsampled to 100 Hz for additional analyses.
-Meanwhile, the Hilbert transform was applied to extract the envelope of each QRS complex, the area under which being taken as a measure of the signal to noise ratio (i.e.
-signal quality index (SQI)).
 For each peak, morphological features (e.g.
@@
 70K).
-Since Mamba is designed for pattern recognition across long contexts, it could potentially interact with signals more effectively than the CRNN if the signals are given directly on a finer time-resolution without windowing the data.
-The CRNN architecture has three input channels for time-series, frequency, and scalar data (see Figure 3): the first two are 3D inputs, with axes for the number of 30-second windows, the window size (i.e.
+Since Mamba is designed for long-context pattern recognition, it could potentially interact with signals more effectively than the CRNN if given them directly without windowing.
+The CRNN architecture has three input channels for time-series, frequency, and scalar data (see Figure 3): the first two are 3D inputs, with axes for the number of 30-second windows (mirroring the 30-second windows used in visual annotation of PSG data according to AASM criteria), the window size (i.e.
 30 seconds worth of sequential data), and number of features, while the 2D scalar inputs have axes for the number of windows and number of features.
@@
 2)
-A recurrent component, consisting of a bidirectional Mamba block that processes the embedding vectors sequentially, followed by an element-wise multi-layer perceptron (MLP) that diminishes the feature-size axis into logits for each class.
-The bidirectional Mamba block (Appendix A) is the concatenation of two Mamba blocks that process epochs in opposite directions.
+A recurrent component, consisting of a bidirectional three-layer Mamba block that processes the embedding vectors sequentially, followed by an element-wise multi-layer perceptron (MLP) that diminishes the feature-size axis into logits for each class.
+The bidirectional Mamba block (Appendix B) is the concatenation of two Mamba blocks that process epochs in opposite directions, allowing for a widened receptive field to detect patterns in signals more effectively.
 The RNN architecture (Figure 2B) processes a single 2D input channel with axes for the number of -second windows and number of features, where is an arbitrary value dividing 30. It comprises:
@@
 2)
-A bidirectional Mamba block that processes windows sequentially.
+A bidirectional three-layer Mamba block that processes windows sequentially.
 3)
@@
 Our previous investigations [30] (Appendix C) have found engineered frequency domain and scalar features to yield significant performance benefits for GRU-based deep learning approaches.
-The engineered scalar features distribute differently depending on sleep stage, while changes to their frequency domains often designate sleep stage transitions (Appendix B).
-Thus, empirical selections of these features (Table 1) were introduced alongside the time-series signals for both the CRNN and RNN.
+The engineered scalar features, which have also shown discriminatory value in other previous models [30], distribute differently depending on sleep stage, while changes to their frequency domains often designate sleep stage transitions (Appendix D).
+Thus, empirical selections of these features (Table 1) were introduced along with the time-series signals for both the CRNN and RNN.
 A series of transformations were applied to further adjust the preprocessed data to the individual architectures.
@@
 Log-magnitude spectrograms were taken with a window size of 100 points and no overlap, then finally concatenated with limb/chest temperature, which was resampled to 1 Hz, and sleep stage label, of which each point was repeated 30 times to 1 Hz.
-RNN CRNN Time-Series Frequency Scalar ECG* Heart Rate (ECG + PPG) ECG* Age PPG* Heart Rate (ECG) PPG* Sex X* Heart Rate (PPG) X* Mean, ECG SQI Y* Pulse Arrival Time (PAT) Y* Mean, PPG SQI Z* Euclidean Norm Minus One (ENMO) Z* Entropy, PPG Chest Temp.
-Z-Angle Z-Angle* Entropy, PPG* Limb Temp.
-Limb-Chest Temperature Difference Skew, PPG* Theta Scale, Heart Rate (ECG + PPG) Phi Kurtosis, ENMO Entropy, Z-Angle
+RNN CRNN Time-Series Frequency Scalar ECG* Heart Rate (ECG + PPG) ECG* Age PPG* Heart Rate (ECG) PPG* Sex X* Heart Rate (PPG) X* Mean, ECG SQI Y* Pulse Arrival Time (PAT) Y* Mean, PPG SQI Z* Euclidean Norm Minus One (ENMO)1 Z* Entropy, PPG Chest Temperature Z-Angle2 Z-Angle* Entropy, PPG* Limb Temperature Limb-Chest Temperature Difference Skew, PPG* Theta Scale, Heart Rate (ECG + PPG) Phi Kurtosis, ENMO Entropy, Z-Angle
 2.4 Training
-The training, validation, and testing sets were randomly split in a size ratio to sizes .
-During the training process, recordings were clipped to random segments to prevent the model from overfitting to the high-level structure of most recordings, which mostly consisted of wake periods at the start and end of the recording with sleep in the middle.
-Cross-entropy loss was weighted on inverse class frequency to adjust for the moderate class imbalance as encapsulated by Figure 1C.
-The model from the iteration that maximized balanced validation accuracy was chosen as the final model.
-Both CRNN and RNN models were trained with the Adam optimizer [32].
-The CRNN model was trained with single batches.
+The training, validation, and testing sets were randomly split in a ratio to sizes .
+During the training process, recordings were clipped to random segments to prevent the model from overfitting to the high-level structure of most recordings, i.e., wake periods at the start and end with sleep in the middle.
+Cross-entropy loss was used, weighted on inverse class frequency to adjust for the moderate class imbalance as encapsulated by Figure 1C.
+The model from the training epoch that maximized balanced validation accuracy was chosen as the final model.
+Both CRNN and RNN models were trained with Adam [32], a widely-used adaptive learning rate optimizer.
+The CRNN model was trained on single batches.
 Recordings were clipped to segments of size varying uniformly between and , and age/sex values individually were set to with probability , an arbitrary value, as a form of data augmentation.
@@
 A visual representation of the ensembling pipeline is presented in Figure 2C.
-As decision tree-based architectures do not output class-wise logits, epoch-wise predictions on the RNN model were generated by aggregating 30 one-second predictions and reducing with the mode.
+2.6 Statistical Analyses
+Results were compiled of 3-class, 4-class, and 5-class sleep staging for the RNN and CRNN architectures and for regular and ensemble models.
+Overall performance was evaluated by aggregating all test set recordings and computing the balanced accuracy, weighted precision, weighted recall, weighted F1, Cohen’s Kappa (), and Matthews correlation coefficient (MCC).
+For the best model type, performance was plotted on a healthy subset of the test set and ablation studies are carried out using various subsets of the ANNE One sensor data and a different sequential architecture.
+For the best 5-class model, Blant-Altman plots were created to evaluate class-wise biases, followed by analyses of averaged per-recording metrics both individually and across clinical variables including age, sex, PLMI, AHI, BMI, TST, and sleep efficiency.
+Because individual recordings do not necessarily contain instances of all five sleep stages, balanced accuracy may not be a sensible metric for recording-wise comparisons; hence, accuracy was presented in lieu of balanced accuracy for those cases.
 3 Results
-Results are compiled of 3-class, 4-class, and 5-class sleep staging for the RNN and CRNN architectures and for regular and ensemble models.
-Overall performance is evaluated by aggregating all test set recordings and computing the balanced accuracy (B.Acc.), precision (Prec.), recall (Rec.), F1, Cohen’s Kappa (), and Matthews correlation coefficient (MCC).
-We additionally conduct analyses of averaged per-recording metrics, biases in the amount of each sleep stage, and differences in performance across age, PLMI, AHI, BMI, TST, and sleep efficiency.
-Because individual recordings do not necessarily contain instances of all five sleep stages, balanced accuracy may not be a sensible metric for recording-wise comparisons; hence, accuracy (Acc.) is presented in lieu of balanced accuracy for those cases.
 3.1 Model Evaluation
@@
 Figure 5 displays the confusion matrices for all -class ensemble models and overall metrics across all -class regular and ensemble models.
-No significant deviations in performance were observed in the RNN models as tested on 1-second windows and as tested on 30-second windows, of which the latter is presented in this paper.
-The RNN ensemble model yields the best metrics and dominates most class-wise metrics (Appendix D).
+No meaningful deviations in performance were observed in the RNN models as tested on 1-second windows and as tested on 30-second windows (balanced accuracies of 1-second RNN ensembles are 83.74%, 75.06%, 64.86% for 3/4/5-class models), of which the latter is presented in this paper.
+The RNN ensemble model yields the best metrics and dominates most class-wise metrics (5-class balanced accuracies of 86.19%, 69.62%, 67.72%, 81.16%, and 86.60% for wake, N1, N2, N3, and REM sleep respectively, Appendix F).
 Ensemble models perform better than their single-model counterparts, while RNNs perform better than the CRNNs despite the shallower feature set and smaller parameter count.
 Most misclassifications are made between N1 and every other class, as well as between N2 and N3 for 5-class sleep staging.
-Lastly, in a set of young and healthy sleepers (age < 40, AHI < 5, PLMI < 5), the RNN ensemble model’s performance is significantly boosted as denoted in the confusion matrices and metrics of Figure 6.
-![[Uncaptioned image]](extracted/6084238/figures/ensemble_outputs.png)
 […]
-![[Uncaptioned image]](extracted/6084238/figures/healthy-subset.png)
-Metric (%) Classes Balanced Accuracy Precision Recall F1 Cohen’s MCC 5 69.26 76.85 70.07 71.55 60.48 61.60 4 79.39 78.82 77.00 76.89 67.23 67.92 3 86.32 87.63 86.70 86.88 77.87 78.14
+In a set of young and healthy sleepers (age < 40, AHI < 5, PLMI < 5), the RNN ensemble model experiences a significant boost in performance as denoted in Figure 6 although removal of these participants from the test set does not substantially worsen model performance, indicating that overall model performance is not driven by these healthy participants (Appendix G, Figure 12).
+Metric (%) Classes Balanced Accuracy Precision Recall F1 Cohen’s MCC 5 69.41 75.30 68.73 69.58 59.83 61.17 4 79.89 78.40 75.83 75.57 66.42 67.58 3 87.67 88.18 87.69 87.78 79.09 79.29
+3.2 Feature Ablation Study
+We next investigated which sets of sensors (chest module vs.
+limb module; ECG vs.
+PPG.
+vs.
+other) were most influential to model performance.
+To do so, modified instances of the RNN ensemble model were trained to determine the utility of its components.
+Firstly, feature sets solely from the ANNE One’s chest module (ECG, accelerometry, chest temperature) or limb module (PPG, limb temperature) were examined.
+As shown in Table 2, the chest-only model trails closely behind the original in performance while the limb-only model exhibits a marked degradation () for all metrics.
+This strongly suggests that the chest module is the principal driver behind model accuracy, although the limb module alone is capable of recovering sleep staging information to a rough degree.
+Next, because the ECG and and PPG sensors themselves are the most likely to experience poor signal quality, we examined the performance of RNN ensemble models trained on only the accelerometry and temperature sensors.
+Remarkably, exclusion of ECG and PPG data resulted in minimal loss of accuracy, meaning that information in the ECG and PPG is also being captured effectively by the accelerometry signals.
+[…]
+3.3 Recurrent Architectures
+In order to understand whether the incorporation of Mamba improved the RNN ensemble model, a similarly-sized variant model replacing Mamba with LSTM was trained, its results provided by Table 3.
+The LSTM model shows a minor loss in all metrics, confirming the Mamba approach is beneficial.
+[…]
 3.4 Global Sleep Metrics
@@
 PSG actual values for a number of summary metrics including time asleep, % N1 sleep, % N2 sleep, % N3 sleep, and % REM sleep.
-The model trends towards slightly underestimating N2 sleep and slightly overestimating N1 sleep and wake.
+The model trends towards slightly underestimating wake and N2 sleep and slightly overestimating N1 and N3 sleep.
 3.5 Performance by Clinical Group
-Averaged recording-wise metrics are cataloged in Table 2 for the 5-class ensemble CRNN and RNN, wherein linear regressions were performed to fit various recording metrics to accuracy.
-Performance of the RNN ensemble is robust to differences in sex, periodic limb movement index, body mass index, total sleep time and sleep efficiency, but shows slightly worse performance with greater age and higher AHI.
-The CRNN model is robust to differences in age, sex, PLMI, and BMI but shows slightly worse performance with higher AHI, shorter TST, and lower sleep efficiency.
-Notably, even where performance varies by clinical factors, the percentage of explained change in performance is nevertheless weak for all metrics ().
-There is no significant difference in accuracy between the raters in the testing set (RNN , CRNN , one-way F-test).
-| RNN | CRNN | ||||||
-| Metric (%) | |||||||
-| Accuracy | 65.09 | 9.84 | 62.58 | 12.50 | |||
-| F1 | 67.69 | 9.92 | 63.81 | 14.17 | |||
-| Cohen’s | 50.39 | 14.91 | 46.33 | 16.13 | |||
-| Matthews Correlation Coefficient | 52.61 | 13.94 | 48.77 | 15.91 | |||
-| Sleep Statistic | |||||||
-| Age | 0.0827 | 0.0350 | 0.0346 | 0.1779 | |||
-| Sex | 0.0689 | 0.0552 | 0.024 | 0.2630 | |||
-| Periodic Limb Movement Index (PLMI) | 0.0120 | 0.4300 | 0.0012 | 0.8061 | |||
-| Apnea-Hypopnea Index (AHI) | 0.1525 | 0.0350 | 0.1622 | 0.0025 | |||
-| Body Mass Index (BMI) | 0.0164 | 0.3559 | 0.0000 | 0.9824 | |||
-| Total Sleep Time (TST) | 0.0000 | 0.9921 | 0.0853 | 0.0322 | |||
-| Sleep Efficiency | 0.0045 | 0.6316 | 0.1503 | 0.0038 | |||
+Averaged recording-wise metrics are catalogued in Table 4 for the 5-class RNN ensemble, wherein linear regressions were performed to fit various recording metrics to accuracy (plots in Appendix G, Figure 13).
+Performance of the RNN ensemble is robust to differences in all metrics except AHI, which had a weak negative correlation with accuracy, though the proportion of variance in accuracy explained by AHI is modest (RNN ).
+Four annotators were involved in generating the ground truth labels used in this study and the agreement between the RNN model and the ground truth labels did not differ by the individual who generated the annotations (RNN , one-way F-test).
+Metric (%) | ||||
+Accuracy | 64.54 | 10.77 | ||
+F1 | 66.72 | 10.28 | ||
+Cohen’s | 50.58 | 14.50 | ||
+Matthews Correlation Coefficient | 52.92 | 13.40 | ||
+Sleep Statistic | ||||
+Age | 0.0618 | 0.0700 | ||
+Sex | 0.0334 | 0.1861 | ||
+Periodic Limb Movement Index (PLMI) | 0.0338 | 0.1830 | ||
+Apnea-Hypopnea Index (AHI) | 0.0751 | 0.0449 | ||
+Body Mass Index (BMI) | 0.0350 | 0.1753 | ||
+Total Sleep Time (TST) | 0.0055 | 0.5940 | ||
+Sleep Efficiency | 0.0002 | 0.9281 | ||
 4 Discussion
 4.1 Key Findings
-This work has successfully developed Mamba-based CRNN and RNN models to accurately perform inference from the ANNE One device for major sleep stages.
-The models demonstrated good performance across a wide range of patient characteristics, without a need to screen and exclude recordings and epochs on the basis of signal quality, and with the RNN architecture slightly outperforming the CRNN architecture.
+This work has developed Mamba-based CRNN and RNN models to accurately perform inference from the ANNE One device for major sleep stages.
+The models demonstrate good performance across a wide range of patient characteristics, without a need to screen and exclude recordings and epochs on the basis of signal quality, and with the RNN architecture slightly outperforming the CRNN architecture.
 4.2 Related Studies
-Previous studies have undertaken the challenge of inferring sleep stage from non-EEG signals (Tables 3 & 4) using a mixture of cardiopulmonary (ECG, PPG, respiratory impedance bands, respiratory flow cannulas) and accelerometry data.
+Previous studies have undertaken the challenge of inferring sleep stage from non-EEG signals (Tables 5 & 7) using a mixture of cardiopulmonary (ECG, PPG, respiratory impedance bands, respiratory flow cannulas) and accelerometry data.
 Many studies used data from conventional PSG recordings [20, 21, 22, 23, 24, 25, 26] although a handful of studies used data from purpose-designed wearable sensors [28, 21, 19, 23, 24, 27].
@@
 Our model exceeded the overall performance of most wearable sensor based models [27, 23, 28] with the exception of one [24].
-Of note, this study examined a younger set of participants (mean age 45) who were free of psychiatric and neurological co-morbidity.
-In contrast, we studied a much older (mean age 57.49) participants with significant sleep co-morbidity (45.81% with AHI > 5).
+Of note, this study examined a younger set of participants (mean age 45) who were free of psychiatric and neurological co-morbidity, which may influence the relationship between sleep stage and its autonomic, cardiac, and pulmonary manifestations.
+In contrast, we studied an older (mean age 57.49) set of participants with significant sleep co-morbidity (45.81% with AHI > 5).
 When we limit our analyses to younger individuals without sleep apnea or significant periodic limb movements, model performance improves further (Figure 6).
@@
 In contrast, in the interest of evaluating real-world performance, we did not exclude any recording or any epochs on the basis of signal quality and the reported results reflect consideration of all available epochs of recording, irrespective of signal quality.
-The robustness of our model to missing data or poor signal quality may be related in part to its use of multiple streams of data such that sensors returning poor quality data are compensated for in the model by those returning good quality data.
+The robustness of our model to missing data or poor signal quality is likely related in part to its use of multiple data streams, in particular accelerometry, which appeared to capture much of the information that would otherwise have been captured by ECG or PPG.
 Of particular comparative interest is the one previous paper addressing sleep stage inference from ANNE One data [19].
-Ultimately, success in developing a population model was limited, and while there was some success in generating person-specific models, the performance of our ensemble RNN model exceeded even the person-specific models generated in this paper.
+Ultimately, success in developing a population model was limited, and while there was some success in generating person-specific models, the performance of our ensemble RNN model exceeded even the best person-specific model in the paper which had a 4-class balanced accuracy of 73.4%.
 […]
@@
 4.3 Model Performance
-Both model architectures obtain maximal per-class performances for wake and REM, and minimal per-class performances for N1 and N2 sleep.
+Both CRNN and RNN model architectures obtain maximal per-class performances for wake and REM, and minimal per-class performances for N1 and N2 sleep.
 As a transitory state, periods of N1 sleep are short and instable, making its discrimination a difficult task.
 Physiological similarities between N2 and N3 sleep, such as cardiorespiratory signatures, similarly complicate the model’s ability to differentiate them, as evidenced by N2-N3 confusion as the most frequent misprediction for 5-class models (as well as N1/N2-N3 confusion for 4-class models).
+Ultimately, the differences betwen N1, N2, and N3 represent a matter of depth rather than kind, and some degree of difficulty in distinguishing these without EEG is to be expected.
 Conversely, wake and REM sleep have clearer differentiating characteristics, such as higher heart rate, which allows for easier classification.
@@
 As well, during epochs where sleep is transitioning between classes, the provided label becomes less reliable; the CRNN attempts to predict a behaviourally inconsistent epoch while it is likely the RNN more flexibly captures this ambiguity by means of multiple differing fine-resolution class predictions throughout the 30 seconds.
-No statistically significant variations in performance based on PLMI, BMI, TST, or sleep efficiency were exhibited by the RNN 5-class ensemble model, providing an important advantage when applied in clinical contexts.
-The significant () association discovered between AHI and accuracy may owe to the fact that individuals with sleep apnea experience sleep fragmentation, resulting in more time in N1 sleep, one of the most-difficult stages to separate; even so, the association was modest ().
-Model performance also correlates between age and accuracy (), though this may in part be explained by the weak agreement between AHI and age ().
+No statistically significant variations in accuracy based on age, sex, PLMI, BMI, TST, or sleep efficiency were exhibited by the RNN 5-class ensemble model, providing an important advantage when applied in clinical contexts.
+The marginally significant () association between AHI and accuracy may owe to the fact that individuals with sleep apnea experience sleep fragmentation, resulting in N1 sleep, one of the most difficult stages to separate; nonetheless, the association was modest ().
+The series of ablation analyses establish that much of the RNN model’s performance was likely driven by chest-module features (i.e.
+ECG, temperature, and xyz accelerometry) rather than limb module (i.e.
+PPG and temperature), meaning that omitting the limb module may be reasonable if participant burden is a concern and collection of oximetry data is not necessary.
+Remarkably, exclusion of both ECG and PPG data with retention of only accelerometry and temperature data resulted in model performance almost indistinguishable from use of all the chest module sensors, suggesting that the inclusion of traditional heart rate sensors (PPG, ECG) is not necessary.
+This highlights the richness of accelerometry data obtained at the chest, spectrograms of which may capture both cardiac (ballistocardiogram) and respiratory data, as well as measures of movement.
+It may also contribute to the robustness of our model given the relative reliability of the accelerometers compared to the ECG/PPG sensors.
 4.4 Limitations, Strengths, and Future Directions
-A few methodological limitations are worth considering.
-1) The differing nature of the way our CRNN and RNN models handle preprocessing and predictions (notably, the 30-second window employed by the CRNN as opposed to the 1-second window used in the RNN) make the models difficult to compare; our reconciling of the RNN to make predictions for 30-second windows may not reflect its full performance.
-2) As neural networks, CRNNs and RNNs provide limited insight into the decoding of sleep stages from the given sensor data.
+This study had many strengths: 1) The ANNE One system itself, with multiple channels of data recorded from a pair or minimally intrusive sensors, is a strength insofar as it provides an element of redundancy should one sensor return low quality signal, 2) the inclusion of Mamba allowed the achievement of relatively high performance with a relatively parameter-scarce architecture, and 3) the study population was varied, including participants from across a spectrum of sleep disorders.
+A few methodological limitations are also worth considering.
+1) The differing nature of the CRNN and RNN in preprocessing and predicting (notably, the CRNN’s 30-second window as opposed to the RNN’s 1-second window) make the models difficult to compare; our reconciling of the RNN to make predictions for 30-second windows may not reflect its full performance.
+2) The black box nature of CRNNs and RNNs make it difficult to ascertain the specific mechanisms used to infer sleep stage from the sensor data.
 Further analyses targeted toward interpreting the neural networks may be explored.
@@
 It is difficult to know for sure how the model will perform in detecting daytime naps during long ambulatory recordings and further studies with ambulatory recordings in clinical populations will be informative.
-This study also had many strengths: 1) The ANNE One system itself, with multiple channels of data recorded from a pair or minimally intrusive sensors, is a strength insofar as it provides an element of redundancy should one sensor return low quality signal, 2) the combination of Mamba with an RNN architecture allowed the achievement of relatively high performance with a relatively parameter-scarce architecture, and 3) the study population was varied, including participants from across a spectrum of sleep disorders.
+4) The recordings were all obtained from a single site.
+Additional work is needed to assess the extent to which these results will be generalizable to ANNE recordings made in different settings on different patient populations.
 5 Conclusion
@@
 Dylan Jones from the Department of Physics at the University of Toronto for offering various forms of support for this research during its infancy.
-Appendix A: Bidirectional Mamba Block
-Appendix B: Scalar and Frequency Features
+Appendix A: Summary Statistics by Groups
+Variable Age 65 Age 65 p AHI AHI 5 p Age [y] 64.01 ± 14.34 51.82 ± 18.07 1.48 Sex [% Female] 44.52 ± 49.70 52.83 ± 49.92 0.123 40.36 ± 49.06 57.29 ± 49.47 1.34 AHI 14.23 ± 16.64 9.90 ± 17.59 0.0203 TST [h] 4.61 ± 1.29 5.43 ± 1.48 1.46 4.86 ± 1.35 5.30 ± 1.52 4.82 PLMI 23.34 ± 35.41 9.22 ± 16.82 7.96 16.64 ± 25.38 13.55 ± 28.18 0.281 Body Mass Index 27.89 ± 4.59 27.90 ± 6.19 0.986 29.45 ± 5.72 26.55 ± 5.13 7.54 Sleep Efficiency [%] 65.81 ± 16.40 75.84 ± 17.81 1.27 69.61 ± 17.48 73.61 ± 18.13 0.0356 Awake [%] 46.64 ± 15.00 37.10 ± 16.47 5.22 43.01 ± 16.08 39.25 ± 16.78 0.0321 N1 Sleep [%] 12.13 ± 6.64 10.43 ± 5.83 0.0110 13.15 ± 6.97 9.37 ± 4.88 5.41 N2 Sleep [%] 25.87 ± 10.31 30.06 ± 10.46 2.17 26.77 ± 11.03 29.72 ± 10.02 8.59 N3 Sleep [%] 8.20 ± 6.73 12.64 ± 8.48 2.38 9.00 ± 7.43 12.41 ± 8.34 6.36 REM Sleep [%] 7.16 ± 5.20 9.77 ± 6.40 5.96 8.07 ± 5.84 9.25 ± 6.22 0.0679
+Appendix B: Bidirectional Mamba Block
 Appendix C: Explorations in Feature Selection
@@
 3-Class Sleep Staging 2-Class Sleep Detection F1 Precision Accuracy F1 Precision Accuracy Wake Time Series Only Baseline 0.681 0.723 0.644 0.740 0.649 0.860 Engineered Features 0.743 0.723 0.764 0.742 0.703 0.786 Engineered Features & Auxiliary Objective 0.727 0.757 0.700 0.781 0.701 0.880 Ensemble 0.732 0.789 0.683 NREM Sleep Time Series Only Baseline 0.534 0.720 0.425 0.793 0.892 0.714 Engineered Features 0.507 0.754 0.381 0.826 0.859 0.826 Engineered Features & Auxiliary Objective 0.619 0.744 0.530 0.834 0.912 0.769 Ensemble 0.702 0.767 0.647 REM Time Series Only Baseline 0.401 0.265 0.827 Engineered Features 0.428 0.287 0.844 Engineered Features & Auxiliary Objective 0.447 0.314 0.778 Ensemble 0.517 0.381 0.679 Macro Time Series Only Baseline 0.539 0.632 0.555 0.767 0.771 0.770 Engineered Features 0.559 0.663 0.581 0.783 0.784 0.792 Engineered Features & Auxiliary Objective 0.600 0.598 0.624 0.807 0.807 0.811 Ensemble 0.651 0.646 0.679
-Appendix D: Class-Wise Evaluation Metrics
-5-Class Sleep Staging 4-Class Sleep Staging 3-Class Sleep Staging B.Acc.
-Prec.
-Rec.
-F1 MCC B.Acc.
-Prec.
-Rec.
-F1 MCC B.Acc.
-Prec.
-Rec.
-F1 MCC Wake Wake Wake CRNN 83.52 84.10 84.13 84.07 67.46 67.53 84.59 84.80 84.75 84.77 69.01 69.02 84.57 84.70 84.45 84.50 68.57 68.70 RNN 86.00 88.85 87.59 87.29 74.01 75.63 88.00 89.57 89.08 88.94 77.36 78.05 88.67 89.26 89.23 89.18 77.90 78.01 CRNN Ens.
-85.63 86.90 86.64 86.49 72.36 72.83 86.96 87.40 87.41 87.38 74.24 74.28 88.05 88.52 88.52 88.49 76.49 76.55 RNN Ens.
-86.68 88.63 87.95 87.75 74.93 75.88 88.49 89.14 89.10 89.04 77.61 77.74 89.03 89.41 89.42 89.39 78.36 78.39 N1 Sleep N1/N2 Sleep Non-REM Sleep CRNN 66.02 86.62 79.98 82.65 22.47 24.23 62.09 67.98 68.51 65.69 26.58 29.20 79.70 80.20 79.98 79.88 59.68 60.00 RNN 67.77 87.06 77.97 81.43 22.61 25.40 68.42 71.65 72.13 71.34 38.39 39.00 81.68 82.36 82.00 81.89 63.72 64.18 CRNN Ens.
-68.82 87.42 80.47 83.18 25.87 28.12 71.45 74.71 75.07 74.32 44.79 45.56 82.29 82.67 82.52 82.46 64.83 65.05 RNN Ens.
-68.96 87.42 79.35 82.44 24.96 27.61 73.52 76.06 76.40 75.96 48.45 48.83 85.15 85.35 85.30 85.28 70.47 70.56 N2 Sleep CRNN 65.60 74.43 76.10 74.58 34.46 35.48 RNN 64.82 76.06 77.46 74.93 34.63 37.49 CRNN Ens.
-65.29 75.54 77.14 75.00 35.03 37.10 RNN Ens.
-68.12 77.33 78.59 77.02 40.56 42.13 N3 Sleep N3 Sleep CRNN 73.95 90.34 89.41 89.83 44.03 44.26 75.29 89.62 81.09 84.14 31.74 35.77 RNN 82.25 91.85 87.58 89.08 47.45 50.21 82.22 91.58 84.59 86.94 42.11 46.62 CRNN Ens.
-77.66 91.19 89.59 90.25 48.04 48.69 78.35 91.05 88.33 89.39 45.80 47.15 RNN Ens.
-81.20 92.14 90.30 91.01 52.80 53.78 82.18 92.12 89.32 90.35 51.18 52.91 REM Sleep REM Sleep REM Sleep CRNN 72.17 91.05 91.33 91.18 45.86 45.90 73.16 90.10 87.77 88.75 38.54 39.38 78.98 92.24 91.29 91.70 52.43 52.79 RNN 82.42 93.23 92.29 92.67 58.19 58.66 82.85 93.17 91.90 92.40 57.33 58.04 87.46 93.49 88.87 90.37 52.70 56.55 CRNN Ens.
-81.70 92.69 91.06 91.70 54.01 54.92 83.06 92.82 90.54 91.38 53.64 55.13 83.47 92.73 89.69 90.78 51.80 53.90 RNN Ens.
-84.79 94.20 93.63 93.86 64.41 64.68 87.16 94.40 93.22 93.65 64.47 65.33 88.17 94.49 92.94 93.48 61.16 65.42
+Appendix D: Scalar and Frequency Features
 Appendix E: UMAP of CRNN Model Learned Embeddings
+Appendix F: Class-Wise Evaluation Metrics
+5-Class Sleep Staging 4-Class Sleep Staging 3-Class Sleep Staging Regular Ensemble Regular Ensemble Regular Ensemble Metric (%) CRNN RNN CRNN RNN CRNN RNN CRNN RNN CRNN RNN CRNN RNN Wake Wake Wake Balanced Accuracy 83.75 84.93 86.47 86.19 84.13 86.12 87.93 87.87 78.57 88.14 87.97 88.27 Precision 84.61 88.16 87.64 88.44 84.48 88.57 88.34 88.67 80.48 88.70 88.24 88.82 Recall 84.68 87.28 87.64 88.04 83.38 88.07 88.32 88.70 76.26 88.72 88.12 88.85 F1 84.63 86.93 87.54 87.81 83.53 87.82 88.33 88.66 76.40 88.71 88.16 88.83 Cohen’s 67.92 72.53 73.93 74.41 66.37 74.42 75.74 76.32 53.63 76.47 75.45 72.89 MCC 67.95 73.94 74.11 75.17 67.04 75.32 75.74 76.37 56.43 76.48 75.49 76.74 N1 Sleep N1/N2 Sleep Non-REM Sleep Balanced Accuracy 62.71 65.07 67.99 69.62 69.22 67.11 71.58 73.47 72.35 83.46 81.03 85.42 Precision 85.30 85.70 86.74 87.09 73.32 68.97 75.37 75.61 74.26 83.60 81.90 85.50 Recall 82.64 78.13 81.74 79.31 73.05 69.43 75.06 75.81 72.28 83.45 80.99 85.41 F1 83.84 81.14 83.77 82.25 71.73 69.01 74.00 75.36 71.73 83.43 80.87 85.41 Cohen’s 21.74 20.73 27.21 26.46 40.67 34.97 45.39 48.25 44.64 66.90 62.01 70.83 MCC 22.12 22.58 28.61 29.06 42.52 35.19 47.03 48.74 46.54 67.05 62.90 70.92 N2 Sleep Balanced Accuracy 64.85 67.47 65.88 67.72 Precision 73.86 74.95 76.25 76.33 Recall 75.49 76.27 77.22 77.50 F1 73.44 75.03 74.80 75.82 Cohen’s 33.51 37.98 36.71 39.58 MCC 35.28 38.87 39.68 41.29 N3 Sleep N3 Sleep Balanced Accuracy 76.48 78.15 79.40 81.16 77.36 81.50 80.08 82.18 Precision 89.80 90.53 90.86 91.71 90.46 90.98 90.92 91.72 Recall 85.36 87.47 87.60 89.66 88.07 81.92 86.85 88.59 F1 87.03 88.64 88.82 90.45 89.02 84.87 88.32 89.72 Cohen’s 39.73 45.03 46.36 52.35 45.54 38.84 45.47 50.70 MCC 41.84 46.52 48.08 53.44 46.58 44.30 47.77 52.60 REM Sleep REM Sleep REM Sleep Balanced Accuracy 75.29 81.55 82.31 86.60 77.83 77.66 82.99 88.67 70.09 85.57 84.13 89.26 Precision 90.95 93.38 92.70 94.69 91.67 93.31 92.63 94.97 90.33 93.47 92.52 94.87 Recall 90.33 93.24 91.39 94.31 90.93 93.76 90.70 94.26 91.04 91.80 89.04 93.81 F1 90.61 93.31 91.91 94.46 91.26 93.46 91.42 94.52 90.63 92.40 90.25 94.18 Cohen’s 47.59 61.83 56.59 69.06 51.59 60.63 55.18 70.06 43.76 60.09 51.95 68.78 MCC 47.72 61.85 57.27 69.23 51.80 61.06 56.35 70.54 44.03 61.23 54.38 69.61
+Appendix G: Additional Model Performance Plots
+Metric (%) Classes Balanced Accuracy Precision Recall F1 Cohen’s MCC 5 63.56 70.91 63.34 65.30 51.43 52.60 4 73.68 75.40 73.15 73.73 60.11 60.46 3 83.13 84.02 83.13 83.38 71.28 71.41

### v4

@@
-Mamba-based Deep Learning Approaches for Sleep Staging on a Wireless Multimodal Wearable System without Electroencephalography
+Mamba-based Deep Learning Approach for Sleep Staging on a Wireless Multimodal Wearable System without Electroencephalography
 Abstract
 Study Objectives
-We investigate Mamba-based deep learning approaches for sleep staging on signals from ANNE One (Sibel Health, Evanston, IL), a non-intrusive dual-module wireless wearable system measuring chest electrocardiography (ECG), accelerometry, and temperature, and finger photoplethysmography (PPG) and temperature.
+We investigate a Mamba-based deep learning approach for sleep staging on signals from ANNE One (Sibel Health, Evanston, IL), a non-intrusive dual-module wireless wearable system measuring chest electrocardiography (ECG), triaxial accelerometry, and chest temperature, and finger photoplethysmography and finger temperature.
 Methods
-We obtained wearable sensor recordings from 360 adults undergoing concurrent polysomnography (PSG) at a tertiary care sleep lab.
+We obtained wearable sensor recordings from 357 adults undergoing concurrent polysomnography (PSG) at a tertiary care sleep lab.
 Each PSG recording was manually scored and these annotations served as ground truth labels for training and evaluation of our models.
 PSG and wearable sensor data were automatically aligned using their ECG channels with manual confirmation by visual inspection.
-We trained Mamba-based models with convolutional-recurrent neural network (CRNN) and the recurrent neural network (RNN) architectures on these recordings.
+We trained a Mamba-based recurrent neural network architecture on these recordings.
 Ensembling of model variants with similar architectures was performed.
 Results
-Our best approach, after ensembling, attains a 3-class (wake, non rapid eye movement [NREM] sleep, rapid eye movement [REM] sleep) balanced accuracy of 84.02%, F1 score of 84.23%, Cohen’s of 72.89%, and a Matthews correlation coefficient (MCC) score of 73.00%; a 4-class (wake, NREM stage 1/2 [N1/N2], NREM stage 3 [N3], REM) balanced accuracy of 75.30%, F1 score of 74.10%, Cohen’s of 61.51%, and MCC score of 61.95%; a 5-class (wake, N1, N2, N3, REM) balanced accuracy of 65.11%, F1 score of 66.15%, Cohen’s of 53.23%, MCC score of 54.38%.
+After ensembling, the model attains a 3-class (wake, non rapid eye movement [NREM] sleep, rapid eye movement [REM] sleep) balanced accuracy of 84.02%, F1 score of 84.23%, Cohen’s of 72.89%, and a Matthews correlation coefficient (MCC) score of 73.00%; a 4-class (wake, light NREM [N1/N2], deep NREM [N3], REM) balanced accuracy of 75.30%, F1 score of 74.10%, Cohen’s of 61.51%, and MCC score of 61.95%; a 5-class (wake, N1, N2, N3, REM) balanced accuracy of 65.11%, F1 score of 66.15%, Cohen’s of 53.23%, MCC score of 54.38%.
 Conclusions
-Deep learning models can infer major sleep stages from the ANNE One, a wearable system without electroencephalography (EEG), and can be successfully applied to data from adults attending a tertiary care sleep clinic.
+Our Mamba-based deep learning model can successfully infer major sleep stages from the ANNE One, a wearable system without electroencephalography (EEG), and can be applied to data from adults attending a tertiary care sleep clinic.
 1University of Toronto Dept. of aMedicine, bMedical Biophysics, cComputer Science Toronto, ON, Canada
 2Sunnybrook Research Institute aDept. of Medicine Neurology Div., bPhysical Sciences Toronto, ON, Canada
-3Vector Institute Toronto, ON, Canada
+3Vector Institute for Artificial Intelligence Toronto, ON, Canada
 4University Health Network Toronto, ON, Canada
@@
 7Queen’s University Dept. of Electrical & Computer Engineering Kingston, ON, Canada
-[…]
+# {andrewhz.zhang, alex.hemo, r.yin}@mail.utoronto.ca
 andrew.lim@utoronto.ca
 Keywords: Deep Learning Sleep Staging Wearable Devices
+Statement of Significance
+Sleep is traditionally measured by in-lab polysomnography followed by visual identification of sleep stages by specialized technologists.
+However, the need for specialized personnel and equipment makes polysomnography difficult to scale, and poses a barrier to many older adults.
+To address this, we developed and evaluated in 357 patients attending a tertiary care sleep clinic a Mamba-based AI model for inferring sleep stage from the ANNE One sensor system – a pair of wearable minimally intrusive sensors placed on the chest and finger.
+We show good performance at distinguishing sleep stages across a range of ages and co-morbid sleep pathologies, suggesting that this may represent a scalable, accurate, minimally intrusive approach to the ambulatory assessment of sleep in clinical populations.
+Graphical Abstract
+![[Uncaptioned image]](figures/graphical_abstract.png)
 1 Introduction
@@
 Sleep is essential to physical and mental health.
-Quantitative evaluation of sleep architecture and physiology is central to the diagnosis of many sleep disorders and plays an important role in evaluating individuals with and at risk for neurodegenerative diseases [1].
-Sleep is traditionally and most reliably measured by polysomnography (PSG), in which brain activity, eye movements, heart rate (HR), muscle tone, and breathing patterns are recorded simultaneously using sensors attached to a patient sleeping in a laboratory environment.
-This is followed by sleep staging, the manual categorization of time periods into different sleep stages based on a trained sleep technician’s visual inspection of the collected physiological data.
-The exact sleep staging methodology depends on the specific framework; under the American Academy of Sleep Medicine’s (AASM) [2] conventions, recordings are divided into 30-second windows, or "epochs", with electroencephalography (EEG), electromyography (EMG), and electro-oculography (EOG) used to classify epochs into three stages: awake, rapid eye movement (REM) sleep, and non-REM (NREM) sleep, which can be further subcategorized as N1, N2, and N3 sleep, with N1 as the lightest and N3 as the deepest NREM sleep.
-EEG is the principal signal used for sleep staging [2, 3, 4].
+Quantitative evaluation of sleep architecture and physiology is central to the diagnosis of many sleep disorders, playing an important role in evaluating individuals with and at risk for neurodegenerative diseases [iranzo2015sleep].
+Sleep is traditionally and most reliably measured by polysomnography (PSG), in which brain activity, eye movements, heart rate (HR), muscle tone, and breathing patterns are recorded simultaneously by sensors attached to a sleeping patient in a laboratory environment.
+This is followed by sleep staging, the categorization of time periods into sleep stages based on a trained sleep technician’s visual inspection of the collected physiological data.
+The exact sleep staging methodology depends on the specific framework; under the American Academy of Sleep Medicine’s (AASM) [aasm] conventions, recordings are divided into 30-second windows, or "epochs".
+Electroencephalography (EEG), electromyography (EMG), and electro-oculography (EOG) are used to classify epochs into three stages: awake, rapid eye movement (REM) sleep, and non-REM (NREM) sleep, which can be further subcategorized as N1, N2, and N3 sleep, with N1 as the lightest and N3 as the deepest NREM sleep.
+EEG is the principal signal used for sleep staging [aasm, rechtschaffen1968brain, imtiaz2021systematic].
 1.2 Wearable Devices and Automated Sleep Staging
 The need for specialized personnel, equipment, and laboratory space makes PSGs expensive to conduct and difficult to scale.
-The cumbersome sensor setup and unfamiliar resting environment may discomfort patients, resulting in sleep that is unrepresentative of that in the home environment.
-Particularly sick or frail older adults may be physically unable to attend in-laboratory PSG, in addition to those living in rural or remote areas without PSG services.
-Furthermore, monitoring patients throughout the night and visually labelling sleep epochs is a time-consuming process that limits the throughput of sleep analyses.
+The cumbersome sensor setup and unfamiliar resting environment may also discomfort patients, resulting in sleep that is unrepresentative of that in the home environment.
+Particularly sick or frail older adults may be physically unable to attend in-laboratory PSG, in addition to those living in remote areas without PSG services.
+Furthermore, monitoring patients throughout the night and manually labelling sleep epochs is a time-inefficient process that limits the volume of sleep analyses.
 In response to these challenges, wearable devices have been developed that record a subset of conventional PSG signals.
-Devices recording EEG have the advantage of directly measuring brain activity but the disadvantage of the discomfort associated with sensors on the head similar to PSGs.
-On the other hand, EEG-less devices are less obtrusive and easily set up for in-home use without a technician, but infer rather than directly measure sleep stage.
+Devices that record EEG have the advantage of directly measuring brain activity but the disadvantage of the discomfort associated with sensors on the head similar to PSGs.
+On the other hand, EEG-less devices are less obtrusive and more easily set up for in-home use without a technician, but infer rather than directly measure sleep stage.
 There has been considerable interest in pairing wearable devices with automated sleep stage inference in order to minimize the need for specialized technologists for labeling and hence improve scalability.
-Indeed, even for the analysis of conventional in-labratory PSG, automated sleep staging has been explored as a means of addressing the time-intensiveness of sleep epoch labelling.
-The focus on pattern recognition and loosely rule-based nature of sleep staging is ideal for machine learning approaches which have employed model architectures like convolutional neural networks (CNNs) [5], recurrent neural networks (RNNs) [6], hybrids of the previous two (e.g.
-convolutional recurrent neural networks (CRNN)) [7], transformers [8], and graph neural networks [9].
-Deep learning models trained on PSG data, most notably the EEG signal or some combination of EEG, EMG, and EOG, have surpassed inter-scorer agreement levels, with 5-class accuracy and Cohen’s as high as 92% and 86% respectively, and can make predictions orders of magnitude faster than a human evaluator at the cost of sacrificed interpretability [10, 11].
+Indeed, even for the analysis of conventional in-laboratory PSG, automated sleep staging has been explored as a means of addressing the time-intensiveness of sleep epoch labelling.
+The focus on pattern recognition and loosely rule-based nature of sleep staging is ideal for machine learning approaches which have employed model architectures like convolutional neural networks (CNNs) [tsinalis2016automatic], recurrent neural networks (RNNs) [malafeev2018automatic], hybrids of the previous two (e.g.
+convolutional recurrent neural networks (CRNN)) [seo2020intra], transformers [phan2022sleeptransformer], and graph neural networks [jia2020graphsleepnet].
+Deep learning models trained on PSG data, most notably the EEG signal or some combination of EEG, EMG, and EOG, have surpassed inter-scorer agreement levels, with 5-class accuracy and Cohen’s as high as 92% and 86% respectively, and can make predictions orders of magnitude faster than a human evaluator at the cost of sacrificed interpretability [fiorillo2019automated, phan2022automatic].
 1.3 Recurrent Neural Networks and Mamba
 RNNs are a class of neural networks that takes in a sequence of vectors (e.g.
-a time series of pulse waves, or breathing) and produce a vector output per input, allowing them to execute a wide breadth of time-series tasks.
-Prevalent modern RNN architectures include long short term memory (LSTM) [12] and gated recurrent units (GRU) [13], which are known to forget very long-range dependencies and whose architecture cannot take advantage of GPU acceleration, limiting the training speed.
-Mamba [14] provides a solution to both problems as a type of state space model (SSM) [15] that combines the excellent long-range dependency modeling of SSMs with an additional time-dependent selection mechanism that allows for focusing on or ignoring particular features, as well as a hardware-friendly algorithm that provides computational efficiency.
-Empirically, Mamba has achieved state-of-the-art performance on supervised and unsupervised metrics using datasets of text, audio waveforms, and DNA sequences [14], while Mamba-based models have been adopted to perform automatic sleep staging using EEG and other PSG signals to high degrees of accuracy [16, 17].
+a time series of pulse waves or breathing) and produce a vector output per input, allowing them to execute a wide breadth of time-series tasks.
+Prevalent modern RNN architectures include long short-term memory (LSTM) [hochreiter1997long] and gated recurrent units (GRU) [cho2014learning], which are known to forget very long-range dependencies and whose architecture cannot exploit GPU acceleration, limiting the training speed.
+Transformers [vaswani2017attention] address these problems but scale poorly in speed for longer sequences both for training and inference.
+Mamba [mamba] as a new generation of RNN provides a solution to these issues as a type of state space model (SSM) [gu2021combining] that combines the excellent long-range dependency modeling of SSMs with a time-dependent selection mechanism that allows for focusing on particular features, as well as a hardware-friendly algorithm that is computationally efficient (see S.1 for details).
+Empirically, Mamba has achieved state-of-the-art performance rivaling that of transformers on supervised and unsupervised metrics from datasets of text, audio waveforms, and DNA sequences [mamba], while Mamba-based models have been adopted to perform sleep staging using EEG and other PSG signals to high degrees of accuracy [zhang2024mssc, zhou2024bit].
+![[Uncaptioned image]](figures/anne_data.drawio.png)
 | Statistic | Range / Count | ||
@@
 1.4 Contribution
-This paper applies Mamba-based deep learning approaches to the problem of sleep staging using data from the ANNE One [18] sensor.
-ANNE One, as seen in Figure 1A, is a flexible, minimally-intrusive, and clinical-grade wireless dual-module wearable system by Sibel Health composed of a finger module, which measures photoplethysmography (PPG) and temperature, and a separate module attached to the chest via an adhesive, which measures electrocardiography (ECG), triaxial accelerometry, and temperature.
-The raw signals of a typical recording are plotted in Figure 1B alongside its hypnogram.
-A prior study centered on ANNE One [19] exists, which primarily emphasized the development of person-specific models which were trained and evaluated on data from the same patient, with a population model trained and evaluated on data from different patients showing much poorer performance.
+This paper applies a Mamba-based deep learning approach to the problem of sleep staging using data from the ANNE One [sibel] sensor.
+ANNE One, as seen in Figure 1A, is a flexible, minimally-intrusive, and clinical-grade wireless dual-module wearable system by Sibel Health.
+It is composed of a finger module, which measures photoplethysmography (PPG) and limb temperature, and a separate module attached to the chest via an adhesive, which measures electrocardiography (ECG), triaxial accelerometry, and chest temperature.
+The raw signals of a typical recording are plotted in Figure 1B alongside a hypnogram, a technician-labelled plot of sleep stage over time.
+A prior study centered on ANNE One [anne_personal] exists, which primarily emphasized the development of person-specific models which were trained and evaluated on data from the same patient; in contrast, the population model, trained and evaluated on data from different patients, showed much poorer performance.
 We aim to develop a superior population model to achieve accurate sleep staging from ANNE One signals recorded from previously unseen patients.
 There is a considerable body of work applying machine learning to the problem of sleep stage inference from non-EEG signals.
-For the most part, these have been trained and tested using the cardiopulmonary sensors of conventional PSG [20, 21, 22, 23, 24, 25, 26], which are typically artifact-free.
-Relatively few studies have applied these approaches to data from wearable devices, which are in general more prone to poor signal quality and signal loss; consequently, such studies often need to reject epochs of poorer-quality signal prior to model training and prediction.
-Moreover, these works have focused on analysis of data from single sensors, typically ECG or PPG [27, 28].
+For the most part, these have been trained and tested using the cardiopulmonary sensors of conventional PSG [radha2019sleep, sun, sleepppg, pini, topalidis2023pulses, jones2024expert, kazemi2024improved], which are typically artifact-free.
+Relatively few studies have applied these approaches to data from wearable devices, which are in general more prone to poor signal quality and signal loss.
+Consequently, such studies often need to reject epochs of poorer-quality signal prior to model training and prediction.
+Moreover, these works have focused on analysis of data from single sensors, typically ECG or PPG [silva, zhang2018sleep].
 Finally, models developed for wearable sensors have largely trained on data from relatively young and healthy volunteers, rather than patients presenting to sleep clinics, raising a question of generalizability to clinical populations.
@@
 This is particularly important as age, sleep, and neurological conditions can all plausibly alter the relationship between sleep stage and autonomic, cardiac, and pulmonary function, meaning that models trained on younger adults may generalize poorly to older adults.
-Finally, this work broadens the Mamba architecture to EEG-less sleep staging by exploring whether Mamba’s relative performance increases [16, 17] are also noticed in this context.
+Finally, this work broadens the Mamba architecture to EEG-less sleep staging by exploring whether Mamba’s relative performance increases [zhang2024mssc, zhou2024bit] are also noticed in this context.
 2 Method
@@
 poor skin contact) overnight.
-Upon completion, each PSG recording was manually labeled by one of four technologists according to AASM standards [2].
+Upon completion, each PSG recording was manually labeled by one of four technologists according to AASM standards [aasm].
 In addition, the portion of the ANNE recordings preceding the onset of PSG recording was labeled as wake based on direct observation of participants by study staff.
@@
 The characteristics of the study’s participants are outlined in Figure 1C.
-Generally, older participants (age 65) and those with sleep apnea had poorer sleep quality, with lower total sleep time, poorer sleep efficiency, and increased N1 sleep time relative to other sleep stages (Appendix A).
-The machine learning task is defined from the dataset as a form of multiclass 1D signal segmentation, wherein the ANNE One signals constitute the features, the PSG-annotated sleep stages form the labels, and class predictions are made per point in time, as demonstrated in Figure 2A.
+Generally, older participants (age 65) and those with sleep apnea had poorer sleep quality, with lower total sleep time, poorer sleep efficiency, and increased N1 sleep time relative to other sleep stages (Figure S.2)
+The machine learning task is defined as multiclass 1D-signal segmentation, wherein the ANNE One signals constitute the features, the PSG-annotated sleep stages form the labels, and class predictions are made per point in time to form a hypnogram as demonstrated in Figure 1B.
 2.2 Preprocessing
-The ECG data were recorded at 512 Hz and preprocessed as follows.
-A wavelet transform was applied to raw ANNE ECG signals to identify temporal windows containing QRS complexes; within each window, raw ECG signals were searched for local maxima and minima to identify the precise timing of the R-peak.
+The ECG data were recorded at 512 Hz and processed by a multilevel 1D Coiflet transform with 5 moments.
+Peaks in the resulting signal were identified and used to define windows; within each window, the original ECG signal was searched for local maxima and minima to determine the precise timing of the R-peak.
 The wavelet-transformed ECG data were subsequently downsampled to 100 Hz for additional analyses.
-For each peak, morphological features (e.g.
-amplitude, sharpness) were extracted and hierarchical density-based spatial clustering of applications with noise (HDBSCAN) [29] served to group all identified peaks into a small number of clusters.
-Peaks occurring in trains with relatively stable RR intervals in the physiological range of human HR were identified as likely representing physiological QRS complexes.
-QRS complexes that clustered morphologically with these peaks, irrespective of regularity, were used to compute beat-to-beat heart rate.
-The PPG data were recorded at 128 Hz and preprocessed as follows.
-First, PPG signals were band-pass filtered between and downsampled to 100 Hz.
-Trains of physiological pulse waves were identified on the basis of relatively stable HR and amplitude.
-These high-quality pulse waves were normalized to a standard amplitude and duration, and clustered using HDBSCAN.
-For each cluster, the mean of its normalized pulse waves was taken as a template wave for that cluster; beat-to-beat template matching was performed to determine the extent to which each pulse wave in the rest of the recording matched one of the template waves, with the lowest sum-squared similarity to the template constituting a measure of signal quality for each PPG pulse wave.
-For each matching pulse wave, the timing of the foot and peak was noted, from which a measure of heart rate was obtained.
-The pulse arrival time (PAT) is the delay from the R-peak of the ECG to the arrival of the pulse wave in the PPG.
-To compute PAT from the ANNE data, the time-series of valid ECG peaks and PPG feet were matched to each other by empirically considering all possible lags within the range msec and selecting that which successfully matched the greatest number of ECG peaks and PPG feet.
-Afterwards, for each ECG R-peak and its corresponding PPG foot, the PAT of that heartbeat was computed, generating a beat-to-beat time series of PAT.
-The chest XYZ accelerometry data were recorded at 210 Hz and low-pass filtered, followed by the computation of the euclidean norm minus one (ENMO), theta (body pitch), phi (body roll) and the z-angle (the deviation from vertical of a vector perpendicular to the face of the sensor).
+We then computed an ECG signal quality index (SQI) as follows.
+Trains of 5 or more physiological QRS complexes were identified on the basis of relatively stable HR (differing by no more than 10 beats per minute between consecutive beats) and amplitude (standard deviation of amplitude ).
+All identified ECG peaks in the recording were then clustered using HDBSCAN [mcinnes2017hdbscan] based on amplitude and morphology; peaks that clustered with the template QRS complexes were considered to have passed quality control (QC).
+Segments of ECG of 2 seconds on either side of these peaks were assigned an SQI of 1 while all other segments were given an SQI of 0.
+The PPG data were recorded at 128 Hz, then band-pass filtered between and downsampled to 100 Hz.
+Trains of 5 or more physiological pulse waves were identified on the basis of relatively stable HR (differing by no more than 10 beats per minute between consecutive beats) and amplitude (differing by no more than 2-fold between beats).
+These high-quality pulse waves were normalized to a standard amplitude and duration, then clustered using HDBSCAN.
+For each cluster, the mean of its normalized pulse waves was taken as a template wave for that cluster.
+Beat-to-beat template matching was then performed to determine the extent to which each pulse wave in the rest of the recording matched one of the template waves, with the lowest sum-squared similarity to the template constituting a measure of signal quality for each PPG pulse wave.
+Pulse waves with minimum sum squared were considered to have passed QC.
+Segments of PPG signal were set to 1 if they were within 2 seconds of a pulse wave passing QC, and 0 otherwise.
+The chest XYZ accelerometry data were recorded at 210 Hz, downsampled to 100 Hz similar to ECG, and band-pass filtered over Hz with a second-order Butterworth filter.
+Log-magnitude spectrograms were taken of each signal by means of a short-time Fourier transform (STFT) with a Hamming window of size equal to the signals’ sampling rate (i.e.
+100) and with no overlap, creating samples at Hz (an arbitrary value that was determined experimentally).
+Only the first 12 frequency bins of the XYZ data were kept as the other higher-frequency bins were made redundant by the band-pass filter.
+Finally, the frequency bins were concatenated with limb/chest temperature, which was recorded at 1 Hz, and sleep stage label, which was created at Hz (as labels are for 30-second windows) and upsampled to 1 Hz by repeating each label 30 times.
 2.3 Model
-Two Mamba-based deep neural network architectures were investigated: a convolutional-recurrent neural network (CRNN) and a recurrent neural network (RNN) as summarized in Figures 2AB, 3.
-The CRNN is a typical sleep stage architecture [21, 11] that emulates the way human evaluators label PSGs: the convolutional component encodes physiological signals within each epoch while the recurrent component captures transitional relationships between sleep stages across epochs.
-The RNN architecture is a lightweight and parameter-scarce alternative to the CRNN, consisting of less parameters (580K vs.
-70K).
-Since Mamba is designed for long-context pattern recognition, it could potentially interact with signals more effectively than the CRNN if given them directly without windowing.
-The CRNN architecture has three input channels for time-series, frequency, and scalar data (see Figure 3): the first two are 3D inputs, with axes for the number of 30-second windows (mirroring the 30-second windows used in visual annotation of PSG data according to AASM criteria), the window size (i.e.
-30 seconds worth of sequential data), and number of features, while the 2D scalar inputs have axes for the number of windows and number of features.
-The model comprises:
+A Mamba-based RNN model (summarized in Figure 2B) was designed to process batches of 2D tensors with axes for the number of -second windows and number of features, which include chest temperature, limb temperature, and frequency bins for ECG, PPG, and XYZ.
+Logits for each sleep stage are returned at each time step.
+A model with output classes has 70K parameters which constitute:
 1)
-A convolutional component, consisting of 1D residual blocks that sequentially downsize the time-series and frequency datas’ window size axes and flatten them into an embedding vector per window.
-The time-series embedding, frequency embedding and scalar data are then concatenated across the feature axis to form an embedding vector per window.
+An element-wise multi-layer perceptron (MLP) with hidden dimensions that diminishes the feature-size axis to dimensions, where and are experimentally determined.
 2)
-A recurrent component, consisting of a bidirectional three-layer Mamba block that processes the embedding vectors sequentially, followed by an element-wise multi-layer perceptron (MLP) that diminishes the feature-size axis into logits for each class.
-The bidirectional Mamba block (Appendix B) is the concatenation of two Mamba blocks that process epochs in opposite directions, allowing for a widened receptive field to detect patterns in signals more effectively.
-The RNN architecture (Figure 2B) processes a single 2D input channel with axes for the number of -second windows and number of features, where is an arbitrary value dividing 30. It comprises:
-1)
-An element-wise MLP with hidden layer dimensions that downsizes the data to dimensions, where the choices of and are arbitrary.
-2)
-A bidirectional three-layer Mamba block that processes windows sequentially.
+A bidirectional three-layer Mamba block.
+This consists of two Mamba blocks [mambapy] that process time windows sequentially in opposite directions, whose outputs are concatenated in the feature-length axis to allow for a widened receptive field that detects patterns in signals more effectively.
 3)
-An element-wise MLP that downsizes the data into logits, which, during inference-time, can be grouped and averaged to form logits per 30-second window to comply with AASM standards.
-Our previous investigations [30] (Appendix C) have found engineered frequency domain and scalar features to yield significant performance benefits for GRU-based deep learning approaches.
-The engineered scalar features, which have also shown discriminatory value in other previous models [30], distribute differently depending on sleep stage, while changes to their frequency domains often designate sleep stage transitions (Appendix D).
-Thus, empirical selections of these features (Table 1) were introduced along with the time-series signals for both the CRNN and RNN.
-A series of transformations were applied to further adjust the preprocessed data to the individual architectures.
-For the CRNN, all signals were resampled to 25 Hz and grouped to 30-second windows, resulting in data points per window.
-Then, depending on feature type (see Table 1),
-For time-series features, the signal was left alone.
-For frequency features, the log-magnitude spectrum was taken of the detrended signal over the 30-second axis.
-For scalar features, the signal was reduced over the 30-second axis.
-Metadata, in the form of patient age and sex (male = 0, female = 1), was incorporated into the scalar features by repeating the constants times throughout the recording.
-For the RNN, the accelerometry, ECG, and PPG signals were detrended, with the accelerometry signals further band-pass filtered over Hz.
-Log-magnitude spectrograms were taken with a window size of 100 points and no overlap, then finally concatenated with limb/chest temperature, which was resampled to 1 Hz, and sleep stage label, of which each point was repeated 30 times to 1 Hz.
-RNN CRNN Time-Series Frequency Scalar ECG* Heart Rate (ECG + PPG) ECG* Age PPG* Heart Rate (ECG) PPG* Sex X* Heart Rate (PPG) X* Mean, ECG SQI Y* Pulse Arrival Time (PAT) Y* Mean, PPG SQI Z* Euclidean Norm Minus One (ENMO)1 Z* Entropy, PPG Chest Temperature Z-Angle2 Z-Angle* Entropy, PPG* Limb Temperature Limb-Chest Temperature Difference Skew, PPG* Theta Scale, Heart Rate (ECG + PPG) Phi Kurtosis, ENMO Entropy, Z-Angle
+An element-wise MLP that downsizes the feature size dimension into logits.
+In accordance with AASM practice, we form 30-second window predictions by grouping 1-second logits into contiguous buckets of 30, taking the mean logit for each class, and outputting the sleep stage with the largest corresponding logit.
 2.4 Training
 The training, validation, and testing sets were randomly split in a ratio to sizes .
-During the training process, recordings were clipped to random segments to prevent the model from overfitting to the high-level structure of most recordings, i.e., wake periods at the start and end with sleep in the middle.
-Cross-entropy loss was used, weighted on inverse class frequency to adjust for the moderate class imbalance as encapsulated by Figure 1C.
-The model from the training epoch that maximized balanced validation accuracy was chosen as the final model.
-Both CRNN and RNN models were trained with Adam [32], a widely-used adaptive learning rate optimizer.
-The CRNN model was trained on single batches.
-Recordings were clipped to segments of size varying uniformly between and , and age/sex values individually were set to with probability , an arbitrary value, as a form of data augmentation.
-The RNN model was trained on a batch size of , with recordings clipped randomly to sequence lengths of seconds.
+During the training process, recordings were clipped to random segments of size 5000 to allow for mini-batch training () and to prevent the model from overfitting to the high-level structure of most recordings, i.e., wake periods at the start and end with sleep in the middle.
+The objective function was cross-entropy loss, weighted on inverse class frequency to adjust for the moderate class imbalance shown in Figure 1C, then averaged across batches and time windows.
+The model was optimized with Adam [kingma], a widely-used adaptive learning rate optimizer, with learning rate 0.001.
+The parameters from the training iteration that maximized balanced validation accuracy were chosen for the final model.
 Three models of varying sleep stage resolution were built: a 3-class, 4-class, and 5-class sleep staging model.
 The 5-class model has classes for wake, N1, N2, N3, and REM sleep, while the 3-class model groups N1, N2, and N3 sleep into non-REM sleep.
-The 4-class model designates N1 and N2 sleep as "light sleep" as opposed to N3 or "deep sleep", a common practice in automatic sleep staging [4].
+The 4-class model designates N1 and N2 sleep as "light sleep" as opposed to N3 or "deep sleep", a common practice in automatic sleep staging [imtiaz2021systematic].
+Separate models were trained for each class count as to treat each class equally.
 2.5 Ensembling
-For each of the classes inside a -class ensemble model, a binary model was trained to predict the presence or absence of that class.
-The logits of these binary models were stacked with logits from the regular -class models to form the inputs to a suite of decision tree-based architectures including AdaBoost [33], random forests [34], extra-tree classifiers [35], histogram gradient boosting [36], and XGBoost [37].
+For each of the classes inside a -class ensemble model, a new binary model of the same architecture was trained from scratch to predict the presence or absence of that class.
+The logits of these binary models were then stacked with logits from the regular 3/4/5-class models.
+This collection of logits formed the inputs to a suite of decision tree-based architectures (see S.3 for details) including AdaBoost [freund1995desicion, scikit-learn], random forests [ho1998random, scikit-learn], extra-tree classifiers [geurts2006extremely, scikit-learn], histogram gradient boosting [cui2021gbht, scikit-learn], and XGBoost [chen2016xgboost].
+This architecture type processes time windows in a time-independent manner, progressively dividing them into smaller and more homogeneous groups by thresholding based on the value of their features (i.e.
+a logit value for a sleep stage); the thresholds and choice of features are learned parameters.
+Sleep stage predictions in a single decision tree are made from the plurality sleep label of training set time windows inside a terminal node, while the proportions of each sleep label become the class probabilities.
 The model among the suite that maximized balanced validation accuracy was selected as the final ensembler model.
@@
 2.6 Statistical Analyses
-Results were compiled of 3-class, 4-class, and 5-class sleep staging for the RNN and CRNN architectures and for regular and ensemble models.
+Results were compiled of 3-class, 4-class, and 5-class sleep staging for the regular and ensemble RNN models.
 Overall performance was evaluated by aggregating all test set recordings and computing the balanced accuracy, weighted precision, weighted recall, weighted F1, Cohen’s kappa (), and Matthews correlation coefficient (MCC).
-For the best model type, performance was plotted on a healthy subset of the test set and ablation studies are carried out using various subsets of the ANNE One sensor data and a different sequential architecture.
-For the best 5-class model, Blant-Altman plots were created to evaluate class-wise biases, followed by analyses of averaged per-recording metrics both individually and across clinical variables including age, sex, PLMI, AHI, BMI, TST, and sleep efficiency.
-Because individual recordings do not necessarily contain instances of all five sleep stages, balanced accuracy may not be a sensible metric for recording-wise comparisons; hence, accuracy was presented in lieu of balanced accuracy for those cases.
+For the best model, performance was plotted on a healthy subset of the test set.
+Ablation studies were carried out using various subsets of the ANNE One sensor data.
+We also retrained the model without downsampling the data, and considered alternative architectures including use of LSTM instead of Mamba, and consideration of a CRNN model utilizing the same preprocessing steps as our RNN model.
+For the best 5-class model, Bland-Altman plots were created to evaluate class-wise biases, followed by analyses of per-recording metrics individually as well as across differing signal quality scenarios, sleep disorders, and clinical variables including age, sex, PLMI, AHI, BMI, TST, and sleep efficiency, and sleep disorders including insomnia, restless legs syndrome, and REM sleep behaviour disorder.
+Because individual recordings did not necessarily contain instances of all five sleep stages, balanced accuracy may not have been a sensible metric for recording-wise comparisons; hence, accuracy was presented in lieu of balanced accuracy for those cases.
 3 Results
 3.1 Model Evaluation
-Typical-case recording predictions for all 5-class models are shown in Figure 4 along with predicted point-wise output class probabilities for models that can produce them.
-A Uniform Manifold Approximation and Projection (UMAP) [38] of an intermediate CRNN output is also provided in Appendix E, which shows the separation of sleep stages as clusters in the learned representation space.
+Typical-case recording predictions for all 5-class models are shown in Figure 3 as hypnograms along with predicted point-wise output class probabilities, known as hypnodensity graphs [anderer].
+A Uniform Manifold Approximation and Projection (UMAP) [umap] of an intermediate model output is also provided in S.4 (Figure S4), which is able to show the separation of sleep stages as clusters in the learned high-dimensional representation space.
 Figure 4 displays the confusion matrices for all -class ensemble models and overall metrics across all -class regular and ensemble models.
-No meaningful deviations in performance were observed in the RNN models as tested on 1-second windows and as tested on 30-second windows (balanced accuracies of 1-second RNN ensembles are 83.74%, 75.06%, 64.86% for 3/4/5-class models), of which the latter is presented in this paper.
-The RNN ensemble model yields the best metrics and dominates most class-wise metrics (5-class balanced accuracies of 86.19%, 69.62%, 67.72%, 81.16%, and 86.60% for wake, N1, N2, N3, and REM sleep respectively, Appendix F).
-Ensemble models perform better than their single-model counterparts, while RNNs perform better than the CRNNs despite the shallower feature set and smaller parameter count.
-Most misclassifications are made between N1 and every other class, as well as between N2 and N3 for 5-class sleep staging.
-[…]
-In a set of young and healthy sleepers (age < 40, AHI < 5, PLMI < 5), the RNN ensemble model experiences a significant boost in performance as denoted in Figure 6 although removal of these participants from the test set does not substantially worsen model performance, indicating that overall model performance is not driven by these healthy participants (Appendix G, Figure 12).
+We plotted performance in both 1-second and 30-second epochs and found similar results irrespective of the epoch length (balanced accuracies of 1-second RNN ensembles are 83.74%, 75.06%, 64.86% for 3/4/5-class models).
+Accordingly, all results in this paper are for 30-second windows.
+Compared to a base RNN model, the ensembling procedure resulted in improved overall performance (5-class balanced accuracies of 86.19%, 69.62%, 67.72%, 81.16%, and 86.60% for wake, N1, N2, N3, and REM sleep respectively, S.5).
+Most misclassifications were made between N1 and every other class, as well as between N2 and N3 for 5-class sleep staging.
+![[Uncaptioned image]](figures/ensemble_outputs_rnn.png)
+Metric (%) Classes Model Balanced Accuracy Precision Recall F1 Cohen’s MCC RNN 60.20 69.48 61.20 63.68 49.16 50.18 5 RNN Ensemble 65.11 71.61 64.41 66.15 53.23 54.38 [1pt/1pt] RNN 67.59 73.08 66.59 68.34 51.95 53.06 4 RNN Ensemble 75.30 75.80 73.68 74.10 61.51 61.95 [1pt/1pt] RNN 81.09 83.20 81.98 82.35 69.59 69.77 3 RNN Ensemble 84.02 84.76 84.04 84.23 72.89 73.00
+In a set of young and healthy sleepers (age < 40, AHI < 5, PLMI < 5), the RNN ensemble model performed even better as denoted in Figure 5 although removal of these participants from the test set did not substantially worsen model performance (S.7, Figure S7), indicating that overall model performance was not driven by this small number of healthy participants.
+![[Uncaptioned image]](figures/healthy-subset.png)
 Metric (%) Classes Balanced Accuracy Precision Recall F1 Cohen’s MCC 5 69.41 75.30 68.73 69.58 59.83 61.17 4 79.89 78.40 75.83 75.57 66.42 67.58 3 87.67 88.18 87.69 87.78 79.09 79.29
@@
 limb module; ECG vs.
-PPG.
-vs.
+PPG vs.
 other) were most influential to model performance.
-To do so, modified instances of the RNN ensemble model were trained to determine the utility of its components.
-Firstly, feature sets solely from the ANNE One’s chest module (ECG, accelerometry, chest temperature) or limb module (PPG, limb temperature) were examined.
-As shown in Table 2, the chest-only model trails closely behind the original in performance while the limb-only model exhibits a marked degradation () for all metrics.
+Modified instances of the ensemble model were trained to determine the utility of its components.
+Firstly, features solely from the ANNE One’s chest module (ECG, accelerometry, chest temperature) or limb module (PPG, limb temperature) were examined.
+As shown in Table 1, the chest-only model trails closely behind the original while the limb-only model exhibits a marked degradation () for all metrics.
 This strongly suggests that the chest module is the principal driver behind model accuracy, although the limb module alone is capable of recovering sleep staging information to a rough degree.
-Next, because the ECG and and PPG sensors themselves are the most likely to experience poor signal quality, we examined the performance of RNN ensemble models trained on only the accelerometry and temperature sensors.
-Remarkably, exclusion of ECG and PPG data resulted in minimal loss of accuracy, meaning that information in the ECG and PPG is also being captured effectively by the accelerometry signals.
-[…]
-3.3 Recurrent Architectures
-In order to understand whether the incorporation of Mamba improved the RNN ensemble model, a similarly-sized variant model replacing Mamba with LSTM was trained, its results provided by Table 3.
-The LSTM model shows a minor loss in all metrics, confirming the Mamba approach is beneficial.
-[…]
+Next, because the ECG and PPG sensors themselves are the most likely to experience poor signal quality, we examined the performance of a model trained on only the accelerometry and temperature sensors.
+Remarkably, exclusion of ECG and PPG data resulted in minimal loss of accuracy, meaning that information in the ECG and PPG is also being captured effectively by the accelerometry signals, with temperature providing additional complementary information.
+Further exclusion of temperature data resulted in an accelerometry-only model that performed similarly to or even slightly better (Table S3), indicating that chest accelerometry alone captures most of the discriminative information used by the network.
+This is corroborated by a per-sensor feature importance analysis with Integrated Gradients [sundararajan2017axiomatic] (Figure S17), which suggests that accelerometry-derived features contribute the most to model performance relative to temperature, PPG, and ECG-derived features.
+Metric (%) Classes Feature Set Balanced Accuracy Precision Recall F1 Cohen’s MCC Full 65.11 71.61 64.41 66.15 53.23 54.38 Chest Module 63.94 70.66 62.63 64.22 51.23 52.59 Limb Module 54.83 62.29 55.08 56.56 41.60 42.59 5 No ECG/PPG 63.85 70.68 63.35 65.03 51.97 53.12 [1pt/1pt] Full 75.30 75.80 73.68 74.10 61.51 61.95 Chest Module 73.46 74.63 71.64 72.26 58.88 59.47 Limb Module 64.89 68.63 63.59 64.66 48.22 49.13 4 No ECG/PPG 73.03 74.46 71.52 72.10 58.70 59.30 [1pt/1pt] Full 84.02 84.76 84.04 84.23 72.89 73.00 Chest Module 81.77 83.58 82.50 82.83 70.39 70.53 Limb Module 73.28 78.10 74.71 75.84 58.21 58.69 3 No ECG/PPG 82.78 84.26 83.18 83.51 71.71 71.56
+To better characterize the information available in chest accelerometry, we probed the signals within these channels.
+At the individual-participant level, aggregated accelerometry spectrograms aligned to the hypnogram (Section S.8) show stage-dependent structure, with clear shifts in power concentrated in a low-frequency band overlapping with the range of expected respiratory rates in sleep.
+This supports the hypothesis that very low-frequency components of chest accelerometry are important for distinguishing sleep stages.
+To examine this quantitatively, we (1) trained a model using accelerometry channels alone and then (2) replaced the band-pass filter with a high-pass cutoff to Hz to attenuate the expected frequency range of resting respirations.
+Results are shown in Table S3.
+The accelerometry-only model performed marginally better than the No ECG/PPG model but worse than the full model; this may be attributable to the data’s unimodal nature, which could reduce the complexity of learning structural patterns in multiple data types during training, thus improving convergence.
+Meanwhile, the reduced performance of the high-passed accelerometry-only model is compatible with the possibility that some information used by the model is in the frequency range of human respirations.
+3.3 Alternative Approaches
+To guard against overfitting and reduce the computation burden, we downsampled the ECG, PPG, and accelerometry signals to 100Hz prior to use in our model.
+To explore the extent to which this would have impacted performance, we retrained the model without downsamping (i.e.
+using 512 Hz ECG, 128 Hz PPG, and 210 Hz accelerometry data) and band-pass filtering.
+This did not result in any decisive improvement in model performance (S.7, Figure S6).
+A strong correlation between high-frequency and low-frequency bins might have contributed to redundancy and increased the likelihood of overfitting to noise.
+In support of this hypothesis, we note the difference in balanced accuracy between the training and test sets was much less in the downsampled data () than in the non-downsampled data ().
+Two variants of the RNN model were produced to determine the value of the Mamba architecture and the 1-second window approach.
+The first, a LSTM-based RNN, replaces the Mamba blocks with LSTM blocks of identical input, hidden, and output dimension sizes, resulting in a comparable parameter count to the original RNN model.
+Following our previous study [sleep_poster], the second is a CRNN (1271K parameters) that groups every 30 time windows into buckets, extracts bucket features with a CNN, and processes features sequentially using a Mamba-based RNN, resulting in predictions made per 30 seconds (see S.6 for details).
+The technique of extracting features from 30 seconds of data for an autoregressive model has seen many variants [sun, sleepppg, zhang2018sleep], making the CRNN suitable as a "baseline" approach to windowing.
+As provided by Table 2 (confusion matrices in S.7, Figure S6), both the LSTM-based RNN models and the CRNN models have modestly reduced performance in all metrics, confirming the Mamba approach and 1-second window approach are beneficial.
+Metric (%) Classes Model Balanced Accuracy Precision Recall F1 Cohen’s MCC CRNN 60.02 56.30 60.02 56.37 46.96 47.92 5 LSTM-Based RNN 62.64 69.46 62.47 64.03 50.76 51.87 RNN 65.11 71.61 64.41 66.15 53.23 54.38 [1pt/1pt] CRNN 70.59 63.68 70.59 66.09 55.13 55.47 4 LSTM-Based RNN 72.82 73.95 71.64 72.06 58.70 59.17 RNN 75.30 75.80 73.68 74.10 61.51 61.95 [1pt/1pt] CRNN 80.21 73.43 80.21 75.76 65.26 65.46 3 LSTM-Based RNN 81.06 82.90 81.37 81.82 68.71 68.95 RNN 84.02 84.76 84.04 84.23 72.89 73.00
 3.4 Global Sleep Metrics
-As a means of evaluating model bias and verifying model robustness, Figure 7 lays out scatter plots and Bland-Altman plots for the 5-class RNN ensemble, comparing RNN predicted vs.
-PSG actual values for a number of summary metrics including time asleep, % N1 sleep, % N2 sleep, % N3 sleep, and % REM sleep.
-The model trends towards slightly underestimating wake and N2 sleep and slightly overestimating N1 and N3 sleep.
-3.5 Performance by Clinical Group
-Averaged recording-wise metrics are catalogued in Table 4 for the 5-class RNN ensemble, wherein linear regressions were performed to fit various recording metrics to accuracy (plots in Appendix G, Figure 13).
-Performance of the RNN ensemble is robust to differences in all metrics except AHI, which had a weak negative correlation with accuracy, though the proportion of variance in accuracy explained by AHI is modest (RNN ).
-Four annotators were involved in generating the ground truth labels used in this study and the agreement between the RNN model and the ground truth labels did not differ by the individual who generated the annotations (RNN , one-way F-test).
+As a means of evaluating model bias and verifying model robustness, Figure 6 contains scatter plots and Bland-Altman plots for the 5-class RNN ensemble, comparing predicted vs.
+actual values for a number of summary metrics including time asleep, % N1 sleep, % N2 sleep, % N3 sleep, and % REM sleep.
+The model trends towards slightly underestimating N2 sleep and wake (as shown by the overestimated sleep in the first column) and slightly overestimating N1 and N3 sleep.
+3.5 Recording-Wise Performance
+We computed recording-wise performance metrics for the 5-class RNN ensemble model, the distributions of which are summarized in Table 3 .
+We then considered recording-level accuracy as a function of clinical and demographic features, and of ECG and PPG signal quality, using linear regression models or -tests (plots in S.7, Figures S8, S9, S11).
+The model performed similarly on recordings from individuals with insomnia as compared to those without (accuracy for those with as compared to without insomnia, ) and in those with RLS compared to those without (accuracy , ).
+Accuracy sunk for those with REM sleep behaviour disorder (accuracy , ), largely driven by poorer classification of REM sleep (39% and 79% of REM sleep was detected by the 5-class RNN ensemble for those with and without RBD respectively).
+Accuracy also declined with higher AHI (); however, this was a modest effect, with accuracy dropping only 2% for every 10 unit higher AHI.
+Accuracy was not significantly associated with age, sex, periodic limb movement index, sleep efficiency, or total sleep time.
+Model accuracy did not vary significantly with the proportion of the recording with bad quality ECG (estimate % decline in accuracy for each 1% increase in the proportion of poor-quality ECG, ) or PPG (estimate % decline in accuracy per 1% increase in the proportion of poor-quality PPG, ), but it was significantly worse in recordings with a high proportion of epochs with neither good ECG nor good PPG SQI (estimate % decline in accuracy for each 1% increase in the proportion of time where both ECG and PPG SQI were poor, ).
+However, this effect was modest () and we note that the proportion of time where this occurred was generally low, with all but 6 recordings having at least one of ECG or PPG with good quality for at least 90% of the recording.
+Even in recordings in which 40% or more of the recording had neither good quality ECG nor good quality PPG, 5-class model accuracy remained above 0.50.
+Four annotators were involved in generating the ground truth labels used in this study and the agreement between the model and the ground truth labels did not differ by the individual who generated the annotations (, one-way F-test).
 | Metric (%) | ||||
@@
 | Sex | 0.0334 | 0.1861 | ||
-Periodic Limb Movement Index (PLMI) | 0.0338 | 0.1830 | ||
-Apnea-Hypopnea Index (AHI) | 0.0751 | 0.0449 | ||
-Body Mass Index (BMI) | 0.0350 | 0.1753 | ||
-Total Sleep Time (TST) | 0.0055 | 0.5940 | ||
+| Periodic Limb Movement Index | 0.0338 | 0.1830 | ||
+| Apnea-Hypopnea Index | 0.0751 | 0.0449 | ||
+| Body Mass Index | 0.0350 | 0.1753 | ||
+| Total Sleep Time | 0.0055 | 0.5940 | ||
 | Sleep Efficiency | 0.0002 | 0.9281 | ||
+| Disorder | ||||
+| Restless Leg Syndrome (9) | 0.0134 | 0.4044 | ||
+| Insomnia (14) | 0.0135 | 0.4032 | ||
+| REM Sleep Behaviour Disorder (5) | 0.0965 | 0.0222 | ||
+| % of Recording with Good SQI In… | ||||
+| ECG | 0.0623 | 0.0687 | ||
+| PPG | 0.0005 | 0.8753 | ||
+| ECG or PPG | 0.0913 | 0.0264 | ||
+| ECG and PPG | 0.0136 | 0.4017 | ||
 4 Discussion
 4.1 Key Findings
-This work has developed Mamba-based CRNN and RNN models to accurately perform inference from the ANNE One device for major sleep stages.
-The models demonstrate good performance across a wide range of patient characteristics, without a need to screen and exclude recordings and epochs on the basis of signal quality, and with the RNN architecture slightly outperforming the CRNN architecture.
+This work has developed a Mamba-based RNN ensemble model to accurately perform inference from the ANNE One device for major sleep stages.
+The model demonstrates good performance across a wide range of patient characteristics, without a need to screen and exclude recordings and epochs on the basis of signal quality.
 4.2 Related Studies
-Previous studies have undertaken the challenge of inferring sleep stage from non-EEG signals (Tables 5 & 7) using a mixture of cardiopulmonary (ECG, PPG, respiratory impedance bands, respiratory flow cannulas) and accelerometry data.
-Many studies used data from conventional PSG recordings [20, 21, 22, 23, 24, 25, 26] although a handful of studies used data from purpose-designed wearable sensors [28, 21, 19, 23, 24, 27].
+Previous studies have undertaken the challenge of inferring sleep stage from non-EEG signals (Tables 4 & 5) using a mixture of cardiopulmonary (ECG, PPG, respiratory impedance bands, respiratory flow cannulas) and accelerometry data.
+Many studies used data from conventional PSG recordings [radha2019sleep, sun, sleepppg, pini, topalidis2023pulses, jones2024expert, kazemi2024improved] although a handful of studies used data from purpose-designed wearable sensors [zhang2018sleep, sun, anne_personal, pini, topalidis2023pulses, silva, wulterkens:2021].
 In general, models trained on and applied to PSG data performed better than models utilizing wearable sensor data.
-This is not surprising since signal quality is often much higher for attended in-laboratory PSGs where more intrusive sensors can be used, and signal quality is continuously monitored and issues can be addressed by the technologist as they occur.
-Even so, despite using wearable rather than PSG sensor data, the performance of our ensemble RNN model approached [21, 24, 25, 26, 20, 22] or exceeded [23] that of models trained on much larger sets of PSG data.
-Our model exceeded the overall performance of most wearable sensor based models [27, 23, 28] with the exception of one [24].
+This is not surprising since signal quality is often much higher for attended in-laboratory PSGs where more intrusive sensors can be used, signal quality is continuously monitored and issues can be addressed by the technologist as they occur.
+Even so, despite using wearable rather than PSG sensor data, the macro performance of our ensemble RNN model approached [sun, topalidis2023pulses, jones2024expert, radha2019sleep, sleepppg] or exceeded [pini, kazemi2024improved] the performance reported for models trained on much larger sets of PSG data.
+Our model exceeded [silva, pini, anne_personal, sun, zhang2018sleep] or matched [wulterkens:2021, schipper] the benchmarks of most wearable sensor based models with the exception of one [topalidis2023pulses].
 Of note, this study examined a younger set of participants (mean age 45) who were free of psychiatric and neurological co-morbidity, which may influence the relationship between sleep stage and its autonomic, cardiac, and pulmonary manifestations.
@@
 When we limit our analyses to younger individuals without sleep apnea or significant periodic limb movements, model performance improves further (Figure 5).
-Another contrast between [24] and the present study is that [24] performed much stricter quality control for recording quality with 54/136 or nearly 40% of nights of wearable sensor recording excluded from analysis due to poor signal quality.
+Another contrast between [topalidis2023pulses] and the present study is that [topalidis2023pulses] performed much stricter QC for recording quality with 54/136 or nearly 40% of nights of wearable sensor recording excluded from analysis due to poor signal quality.
 In contrast, in the interest of evaluating real-world performance, we did not exclude any recording or any epochs on the basis of signal quality and the reported results reflect consideration of all available epochs of recording, irrespective of signal quality.
-The robustness of our model to missing data or poor signal quality is likely related in part to its use of multiple data streams, in particular accelerometry, which appeared to capture much of the information that would otherwise have been captured by ECG or PPG.
-Of particular comparative interest is the one previous paper addressing sleep stage inference from ANNE One data [19].
+The robustness of our model to missing data or poor signal quality is likely related in part to its use of multiple data streams, in particular accelerometry, which appeared to capture much of the information that would otherwise have been captured by ECG or PPG, such that sensors returning poor-quality data are compensated for in the model by those returning good quality data.
+Indeed, poor signal quality was present in at least one sensor module for 866 (29.7%) of the 2914 hours of recording in our dataset.
+However, for 725 (83.7%) of these hours, the other module returned sufficient quality signal to allow accurate predictions.
+Of particular comparative interest is the one previous paper addressing sleep stage inference from ANNE One data [anne_personal].
 Ultimately, success in developing a population model was limited, and while there was some success in generating person-specific models, the performance of our ensemble RNN model exceeded even the best person-specific model in the paper which had a 4-class balanced accuracy of 73.4%.
-[…]
-[…]
+Study (Year) Dataset Metric % Signal Features Radha et al.
+(2019) [radha2019sleep] SIESTA (7 sleep labratories) 584 recordings (292 subjects) 25% testing 4-stage Acc.
+4-stage 77 61 ECG [1pt/1pt] Sun et al.
+(2020) [sun] MGH (sleep laboratory) 8682 recordings (7208 subjects) 20% testing 3-stage B.Acc.
+3-stage F1 3-stage 5-stage B.Acc.
+5-stage F1 5-stage 86.8 84.2 76.0 70.3 68.1 58.5 ECG, Resp.
+Effort [1pt/1pt] Kotzen et al.
+(2022) [sleepppg] SHHS Visit 1 (multi-center) 5758 recordings (pretraining) MESA (general population) 2056 recordings 204 testing 4-stage Acc.
+4-stage B.Acc.
+4-stage F1 4-stage 84 76 83 75 PPG [1pt/1pt] Kotzen et al.
+(2022) [sleepppg] SHHS Visit 1 (multi-center) 5758 recordings (pretraining) CFS Visit-5 v1 (family study) 324 recordings 80 testing 4-stage Acc.
+4-stage B.Acc.
+4-stage F1 4-stage 82 80 82 74 PPG [1pt/1pt] Pini et al.
+(2022) [pini] Unspecified (academic sleep centers) 12404 recordings CinC (sleep laboratory) 994 recordings External testing set 3-stage B.Acc.
+3-stage F1 3-stage 3-stage MCC 4-stage B.Acc.
+4-stage F1 4-stage 4-stage MCC 74.0 81.4 61.6 61.7 63.6 72.0 53.6 54.0 ECG [1pt/1pt] Topalidis et al.
+(2023) [topalidis2023pulses] Unspecified (general population w/ sleep complaints) 314 recordings (185 subjects) 25% testing Ambulatory home recordings 4-stage B.Acc 4-stage F1 4-stage 4-stage MCC 85.4 86.3 79.2 79.2 ECG [1pt/1pt] Jones et al.
+(2024) [jones2024expert] CCSHS, CFS, CHAT, MESA, WSC 4000 random sampled recordings 500 testing 5-stage B.Acc 5-stage Median 74 72.5 ECG [1pt/1pt] Kazemi et al.
+(2024) [kazemi2024improved] UC Irvine Sleep Center 123 recordings 20% testing 3-stage 4-stage 5-stage 71.4 55.0 61.6 PPG, Resp.
+Flow Rate, Resp.
+Effort
+Study (Year) Dataset Metric % Signal Features Zhang et al.
+(2018) [zhang2018sleep] Beijing General Hospital of the Air Force (general population) 39 recordings 12.5% testing 5-stage Prec.
+5-stage Rec.
+5-stage F1 58.5 61.1 58.5 Heart Rate, Actigraphy [1pt/1pt] Sun et al.
+(2020) [sun] Model from Table 4 SHHS Visit 1 & 2 (multi-center) 1000 random sampled recordings External testing set 3-stage B.Acc.
+3-stage F1 3-stage 5-stage B.Acc.
+5-stage F1 5-stage 80.8 80.2 69.7 63.9 58.6 53.3 ECG, Resp.
+Effort [1pt/1pt] Wulterkens et al.
+(2021) [wulterkens:2021] SOMNIA, N2N, HHS, HealthBed 835 recordings 35% testing 4-stage B.Acc.
+4-stage F1.
+4-stage 4-stage MCC 75.4 76.4 63.7 63.7 PPG, Accel.
+[1pt/1pt] Chen et al.
+(2022) [anne_personal] Individual-Based Model Shirley Ryan AbilityLab (inpatient unit) 10 recordings 10% testing 3-class Mean 4-class B.Acc.
+4-class Mean 60.0 73.0 61.7 ECG, PPG, Accel., Temp.
+[1pt/1pt] Chen et al.
+(2022) [anne_personal] Population Model Shirley Ryan AbilityLab (inpatient unit) 10 recordings 10% testing 3-class Mean 4-class Mean 17.1 6.1 ECG, PPG, Accel., Temp.
+[1pt/1pt] Pini et al.
+(2022) [pini] Model from Table 4 Z3Pulse (general population) 156 recordings (52 subjects) ECG-based wearable External testing set 3-stage B.Acc.
+3-stage F1 3-stage 3-stage MCC 4-stage B.Acc.
+4-stage F1 4-stage 4-stage MCC 72.5 79.8 61.2 61.4 64.2 69.8 53.2 53.6 ECG [1pt/1pt] Topalidis et al.
+(2023) [topalidis2023pulses] Model from Table 4 Unspecified (general population w/ sleep complaints) 314 recordings (185 subjects) 25% testing 4-stage B.Acc 4-stage F1 4-stage 4-stage MCC 84.4 83.9 76.2 76.3 ECG, PPG [1pt/1pt] Schipper et al.
+(2024) [schipper] Sleep Medicine Center Kempenhaeghe, OLVG Hospital (at-home recordings) 323 recordings 4-stage B.Acc.
+4-stage F1 4-stage 4-stage MCC 75.6 79.4 67.1 67.2 Accel.
+[1pt/1pt] Silva et al.
+(2024) [silva] Instituto do Sono (general population) 1522 recordings (1430 subjects) 586 testing 4-stage Mean Acc.
+4-stage Mean 70.7 56 Heart Rate, Accel.
 4.3 Model Performance
-Both CRNN and RNN model architectures obtain maximal per-class performances for wake and REM, and minimal per-class performances for N1 and N2 sleep.
+The RNN model architecture obtained maximal per-class performances for wake and REM, and minimal per-class performances for N1 and N2 sleep.
 As a transitory state, periods of N1 sleep are short and instable, making its discrimination a difficult task.
-Physiological similarities between N2 and N3 sleep, such as cardiorespiratory signatures, similarly complicate the model’s ability to differentiate them, as evidenced by N2-N3 confusion as the most frequent misprediction for 5-class models (as well as N1/N2-N3 confusion for 4-class models).
-Ultimately, the differences betwen N1, N2, and N3 represent a matter of depth rather than kind, and some degree of difficulty in distinguishing these without EEG is to be expected.
+Physiological similarities between N2 and N3 sleep, such as cardiorespiratory signatures, similarly complicate the model’s ability to differentiate them, as evidenced by N2–N3 confusion as the most frequent misprediction for 5-class models (as well as N1/N2–N3 confusion for 4-class models).
+Of note, these classes also represent those hardest to distinguish from each other in sleep staging by human annotators [rosenberg:2013].
 Conversely, wake and REM sleep have clearer differentiating characteristics, such as higher heart rate, which allows for easier classification.
-In spite of a simpler feature set and smaller parameter count, the RNNs outperform the CRNNs in virtually every metric.
-It is possible that the independent processing of segmented windows by the CNN component interfered with the wider CRNN’s ability to discern changes at neighbouring window boundaries.
-As well, during epochs where sleep is transitioning between classes, the provided label becomes less reliable; the CRNN attempts to predict a behaviourally inconsistent epoch while it is likely the RNN more flexibly captures this ambiguity by means of multiple differing fine-resolution class predictions throughout the 30 seconds.
-No statistically significant variations in accuracy based on age, sex, PLMI, BMI, TST, or sleep efficiency were exhibited by the RNN 5-class ensemble model, providing an important advantage when applied in clinical contexts.
+No statistically significant variations in accuracy based on age, sex, PLMI, BMI, TST, sleep efficiency, insomnia, or RLS were exhibited by the 5-class ensemble model, providing an important advantage when applied in clinical contexts.
 The marginally significant () association between AHI and accuracy may owe to the fact that individuals with sleep apnea experience sleep fragmentation, resulting in N1 sleep, one of the most difficult stages to separate; nonetheless, the association was modest ().
-The series of ablation analyses establish that much of the RNN model’s performance was likely driven by chest-module features (i.e.
+The model showed poorer performance in individuals with RBD, driven by difficulties in classifying REM sleep.
+This may reflect in part the centrality of accelerometry to our model, which, in the context of abnormal movements in REM sleep, may be apt to misclassify REM sleep.
+The series of ablation analyses establish that much of the RNN model’s performance was likely fueled by chest module features (i.e.
 ECG, temperature, and XYZ accelerometry) rather than limb module (i.e.
 PPG and temperature), meaning that omitting the limb module may be reasonable if participant burden is a concern and collection of oximetry data is not necessary.
-Remarkably, exclusion of both ECG and PPG data with retention of only accelerometry and temperature data resulted in model performance almost indistinguishable from use of all the chest module sensors, suggesting that the inclusion of traditional heart rate sensors (PPG, ECG) is not necessary.
-This highlights the richness of accelerometry data obtained at the chest, spectrograms of which may capture both cardiac (ballistocardiogram) and respiratory data, as well as measures of movement.
-It may also contribute to the robustness of our model given the relative reliability of the accelerometers compared to the ECG/PPG sensors.
+Exclusion of both ECG and PPG data with retention of only accelerometry and temperature data resulted in model performance almost indistinguishable from use of all the chest module sensors, suggesting that the inclusion of traditional heart rate sensors (PPG, ECG) is not necessary.
+This highlights the richness of accelerometry data obtained at the chest, which may capture respiratory data as well as measures of movement, and may also contribute to the robustness of our model given the relative reliability of the accelerometers compared to the ECG/PPG sensors.
+At the same time, the attenuation of low-frequency activity (S.9) produced only modest degradations in performance, suggesting that the model is exploiting a nuanced combination of low-frequency respiratory motion, postural changes, and other movements, rather than any single narrow frequency band alone.
 4.4 Limitations, Strengths, and Future Directions
-This study had many strengths: 1) The ANNE One system itself, with multiple channels of data recorded from a pair or minimally intrusive sensors, is a strength insofar as it provides an element of redundancy should one sensor return low quality signal, 2) the inclusion of Mamba allowed the achievement of relatively high performance with a relatively parameter-scarce architecture, and 3) the study population was varied, including participants from across a spectrum of sleep disorders.
+The core strengths of this study are: 1) the ANNE One system itself, with multiple channels of data recorded from a pair of minimally intrusive sensors, is a strength insofar as it provides an element of redundancy should one sensor return low quality signal, 2) the inclusion of Mamba allowed the achievement of relatively high performance with a relatively parameter-scarce architecture, and 3) the study population was large and varied, including participants from across a spectrum of sleep disorders.
 A few methodological limitations are also worth considering.
-1) The differing nature of the CRNN and RNN in preprocessing and predicting (notably, the CRNN’s 30-second window as opposed to the RNN’s 1-second window) make the models difficult to compare; our reconciling of the RNN to make predictions for 30-second windows may not reflect its full performance.
-2) The black box nature of CRNNs and RNNs make it difficult to ascertain the specific mechanisms used to infer sleep stage from the sensor data.
-Further analyses targeted toward interpreting the neural networks may be explored.
-3) The recordings were overnight with a brief period of wakefulness before lights out, resulting in a preponderance of sleep vs.
-wake.
-It is difficult to know for sure how the model will perform in detecting daytime naps during long ambulatory recordings and further studies with ambulatory recordings in clinical populations will be informative.
-4) The recordings were all obtained from a single site.
-Additional work is needed to assess the extent to which these results will be generalizable to ANNE recordings made in different settings on different patient populations.
+1) The black box nature of the RNN makes it difficult to ascertain the specific mechanisms used to infer sleep stage from the sensor data.
+2) The recordings were overnight with a brief period of wakefulness before lights out, resulting in a preponderance of sleep vs.
+wake, so it is difficult to know for sure how the model will perform in detecting daytime naps during long ambulatory recordings, and further studies with ambulatory recordings in clinical populations will be informative.
+3) The recordings were all obtained from a single site, and additional work is needed to assess the extent to which these results will be generalizable to ANNE recordings made in different settings on different patient populations.
+The choice of 1-second windows created both advantages and disadvantages.
+It reduced the number of frequency bins per window, saving on model parameter count, and inflated the number of epochs by thirtyfold, which Mamba’s long-term dependencies could effectively capture.
+However, this setup complicated interpretability: while high-frequency content ( Hz) emerged in the model’s frequency bin features, low-frequency content ( Hz) manifested as changes across time windows, making their effects more difficult to capture with feature importance techniques.
+An important future direction will be to better delineate how chest accelerometry contributes so strongly to model robustness, particularly in the absence of ECG and PPG.
+Our results suggest that some of this may be related to low-frequency, possibly respiratory, activity.
+However, the model appears to be making use of other information as well, and defining the physiological basis of this will be critical to better understanding the potential and limitations of chest accelerometry-based sleep staging.
 5 Conclusion
-This work introduces CRNN and RNN models to automatically perform 3-class, 4-class, and 5-class sleep staging using data from the ANNE One dual sensor using sleep stages inferred from concurrent PSG recordings, with the RNN ensemble model topping the others in virtually every performance metric.
-The model meets or surpasses metrics in related literature, especially for wearable devices, and demonstrates robustness across a wide range of clinical characteristics.
+This work introduces a Mamba-based RNN ensemble model to automatically perform 3-class, 4-class, and 5-class sleep staging using data from the ANNE One dual sensor using sleep stages inferred from concurrent PSG recordings.
+The model meets or surpasses performance metrics in related literature when applied to our ANNE One dataset, especially for wearable devices, and demonstrates robustness across a wide range of clinical characteristics.
 Moreover, the model’s capability was achieved without prescreening and excluding large numbers of patients or epochs for reasons of signal quality.
@@
 Our model provides a scalable, accurate minimally-intrusive approach to the ambulatory assessment of sleep staging without the need for EEG.
-Acknowledgments
+Financial Disclosure Statement
+The authors have no financial conflicts of interest to disclose.
 This research is supported by the Centre for Aging and Brain Health Innovation, Canadian Institutes of Health Research, and National Institute on Aging.
 This research is not funded by Sibel Health, the manufacturer of the ANNE One device.
-Special thanks to Prof.
-Jimmy Ba from the Department of Computer Science; Prof.
-Paul Kushner and Prof.
-Dylan Jones from the Department of Physics at the University of Toronto for offering various forms of support for this research during its infancy.
-Appendix A: Summary Statistics by Groups
+Non-Financial Disclosure Statement
+The authors have no non-financial conflicts of interest to disclose.
+Acknowledgments
+Special thanks to Prof. Paul Kushner and Prof. Dylan Jones from the Department of Physics at the University of Toronto for offering various forms of support for this research during its infancy.
+Data Availability
+All data will be made available through the Canadian Federated Research Data Repository at https://www.frdr-dfdr.ca/repo/ upon publication.
+All code will be made available at https://github.com/OSHS2019/ upon publication.
+S Supplementary Materials
+S.1 Mamba Architecture
+S.1.1 Selective SSM
+At the core of the Mamba architecture is the state space model (SSM).
+For some single channel input , -dimensional latent state , single channel output , and continuous time , a SSM is governed by a system of first-order linear ordinary differential equations (ODE):
+| where , , , are matrices that parameterize the SSM.
+However, in the context of sequence modeling, we require the SSM to work with discrete time steps .
+To this end we apply zero-order hold (ZOH) discretization to and in (S1) such that: | ||||
+| where is the ZOH discretization time step, which governs the magnitude of each discrete update to the system’s trajectory.
+As such, we can reformulate the SSM from a system of ODEs to a recurrence: | ||||
+| Note that Mamba’s SSM still learns the parameters and rather than the discretized and directly, where ZOH discretization is applied during forward pass.
+The Mamba architecture also imposes an additional structural constraint that is diagonal.
+The SSM as formulated by (S5) and (S6) is linear time-invariant (LTI), which means that all parameters remain static and input-independent across time steps.
+This makes encoding context-awareness and adaptive behavior challenging.
+Mamba’s selective SSM breaks LTI, building input dependence into (S5) and (S6) by reparameterizing and .
+For a length input sequence : | ||||
+| where is a parameterized linear projection from the input to .
+Mamba’s selective SSM additionally reparameterizes to make the ZOH discretization time step adaptable to the input such that it functions similarly to RNN gating mechanisms, which is when the magnitude of each hidden state update is modulated as a function of the input: | ||||
+where .
+For a multi-channel input with channels , the selective SSM is applied channel-wise with a total of hidden states.
+Notice that by making the selective SSM not LTI, the Mamba architecture’s every hidden state update is dependent on the entire history of input such that it can capture arbitrarily long-range dependencies across the entire sequence similar to the Transformer architecture.
+S.1.2 Mamba Block
+A Mamba block additionally wraps the selective SSM (S.1.1) around linear projections, residual connections, and convolutions.
+The implementation of a Mamba block used in this study [mambapy] is as shown on Figure S1.
+S.1.3 Bidirectional Mamba Block
+Our bidirectional Mamba block consists of two Mamba blocks (S.1.2).
+One models the sequence in chronological order, while the other models the sequence in reverse-chronological order, as shown on Figure S2.
+S.2 Summary Statistics by Groups
 Variable Age 65 Age 65 p AHI AHI 5 p Age [y] 64.01 ± 14.34 51.82 ± 18.07 1.48 Sex [% Female] 44.52 ± 49.70 52.83 ± 49.92 0.123 40.36 ± 49.06 57.29 ± 49.47 1.34 AHI 14.23 ± 16.64 9.90 ± 17.59 0.0203 TST [h] 4.61 ± 1.29 5.43 ± 1.48 1.46 4.86 ± 1.35 5.30 ± 1.52 4.82 PLMI 23.34 ± 35.41 9.22 ± 16.82 7.96 16.64 ± 25.38 13.55 ± 28.18 0.281 Body Mass Index 27.89 ± 4.59 27.90 ± 6.19 0.986 29.45 ± 5.72 26.55 ± 5.13 7.54 Sleep Efficiency [%] 65.81 ± 16.40 75.84 ± 17.81 1.27 69.61 ± 17.48 73.61 ± 18.13 0.0356 Awake [%] 46.64 ± 15.00 37.10 ± 16.47 5.22 43.01 ± 16.08 39.25 ± 16.78 0.0321 N1 Sleep [%] 12.13 ± 6.64 10.43 ± 5.83 0.0110 13.15 ± 6.97 9.37 ± 4.88 5.41 N2 Sleep [%] 25.87 ± 10.31 30.06 ± 10.46 2.17 26.77 ± 11.03 29.72 ± 10.02 8.59 N3 Sleep [%] 8.20 ± 6.73 12.64 ± 8.48 2.38 9.00 ± 7.43 12.41 ± 8.34 6.36 REM Sleep [%] 7.16 ± 5.20 9.77 ± 6.40 5.96 8.07 ± 5.84 9.25 ± 6.22 0.0679
-Appendix B: Bidirectional Mamba Block
-Appendix C: Explorations in Feature Selection
-A previous iteration of the model used a pure CRNN (GRU) approach, and the approach encompassed two additional strategies, which are utilized within the paper without extensive justification.
-The engineered features are a suite of additional features across the time, frequency, and scalar domains used alongside the raw signals as extracted from the ANNE sensors.
-Time Domain Frequency Domain Scalar Domain PPG, ECG Heart Rates PPG Freq.
-Spectra PPG Freq.
-Spectra Skew PPG, ECG Signal Quality Index ECG Freq.
-Spectra PPG Time Entropy PPG, ECG Amplitude x-accel.
-Freq.
-Spectra PPG Freq.
-Spectra Entropy Pulse Arrival Time (PAT) y-accel.
-Freq.
-Spectra | Heart Rate| Accel.
-ENMO z-accel.
-Freq.
-Spectra Accel.
-ENMO Kurtosis SpO2 Accel.
-z-angle Freq.
-Spectra Entropy Metadata (Gender, Age)
-In addition, the auxiliary training objective was introduced in the form of the transition state.
-These states are derived directly from the target labels.
-With this, the model attempts to predict whether there will be a change in sleep class from this epoch to the next.
-The rationale is that sleep stages in reality do not always change at the end of 30-second windows, and many epochs may have data that is transitioning between two separate sleep stages, which would be difficult for the model as they may not resemble other members of their labelled class and instead as a mixture of both sleep stages.
-By training the model to also predict correct transitions, it should assist the model in determining which epochs may be occurring during a transition between stages, and thus make it more robust in general.
-The application of the transition state as an auxiliary objective was attempted in the Mamba approaches in the main paper; however, it did not yield a noticeable increase in performance, and thus was not included in the final models.
-The results of the approach are detailed in the graph below, with results indicating that the two approaches mentioned positively impacted results.
-3-Class Sleep Staging 2-Class Sleep Detection F1 Precision Accuracy F1 Precision Accuracy Wake Time Series Only Baseline 0.681 0.723 0.644 0.740 0.649 0.860 Engineered Features 0.743 0.723 0.764 0.742 0.703 0.786 Engineered Features & Auxiliary Objective 0.727 0.757 0.700 0.781 0.701 0.880 Ensemble 0.732 0.789 0.683 NREM Sleep Time Series Only Baseline 0.534 0.720 0.425 0.793 0.892 0.714 Engineered Features 0.507 0.754 0.381 0.826 0.859 0.826 Engineered Features & Auxiliary Objective 0.619 0.744 0.530 0.834 0.912 0.769 Ensemble 0.702 0.767 0.647 REM Time Series Only Baseline 0.401 0.265 0.827 Engineered Features 0.428 0.287 0.844 Engineered Features & Auxiliary Objective 0.447 0.314 0.778 Ensemble 0.517 0.381 0.679 Macro Time Series Only Baseline 0.539 0.632 0.555 0.767 0.771 0.770 Engineered Features 0.559 0.663 0.581 0.783 0.784 0.792 Engineered Features & Auxiliary Objective 0.600 0.598 0.624 0.807 0.807 0.811 Ensemble 0.651 0.646 0.679
-Appendix D: Scalar and Frequency Features
-Appendix E: UMAP of CRNN Model Learned Embeddings
-Appendix F: Class-Wise Evaluation Metrics
-5-Class Sleep Staging 4-Class Sleep Staging 3-Class Sleep Staging Regular Ensemble Regular Ensemble Regular Ensemble Metric (%) CRNN RNN CRNN RNN CRNN RNN CRNN RNN CRNN RNN CRNN RNN Wake Wake Wake Balanced Accuracy 83.75 84.93 86.47 86.19 84.13 86.12 87.93 87.87 78.57 88.14 87.97 88.27 Precision 84.61 88.16 87.64 88.44 84.48 88.57 88.34 88.67 80.48 88.70 88.24 88.82 Recall 84.68 87.28 87.64 88.04 83.38 88.07 88.32 88.70 76.26 88.72 88.12 88.85 F1 84.63 86.93 87.54 87.81 83.53 87.82 88.33 88.66 76.40 88.71 88.16 88.83 Cohen’s 67.92 72.53 73.93 74.41 66.37 74.42 75.74 76.32 53.63 76.47 75.45 72.89 MCC 67.95 73.94 74.11 75.17 67.04 75.32 75.74 76.37 56.43 76.48 75.49 76.74 N1 Sleep N1/N2 Sleep Non-REM Sleep Balanced Accuracy 62.71 65.07 67.99 69.62 69.22 67.11 71.58 73.47 72.35 83.46 81.03 85.42 Precision 85.30 85.70 86.74 87.09 73.32 68.97 75.37 75.61 74.26 83.60 81.90 85.50 Recall 82.64 78.13 81.74 79.31 73.05 69.43 75.06 75.81 72.28 83.45 80.99 85.41 F1 83.84 81.14 83.77 82.25 71.73 69.01 74.00 75.36 71.73 83.43 80.87 85.41 Cohen’s 21.74 20.73 27.21 26.46 40.67 34.97 45.39 48.25 44.64 66.90 62.01 70.83 MCC 22.12 22.58 28.61 29.06 42.52 35.19 47.03 48.74 46.54 67.05 62.90 70.92 N2 Sleep Balanced Accuracy 64.85 67.47 65.88 67.72 Precision 73.86 74.95 76.25 76.33 Recall 75.49 76.27 77.22 77.50 F1 73.44 75.03 74.80 75.82 Cohen’s 33.51 37.98 36.71 39.58 MCC 35.28 38.87 39.68 41.29 N3 Sleep N3 Sleep Balanced Accuracy 76.48 78.15 79.40 81.16 77.36 81.50 80.08 82.18 Precision 89.80 90.53 90.86 91.71 90.46 90.98 90.92 91.72 Recall 85.36 87.47 87.60 89.66 88.07 81.92 86.85 88.59 F1 87.03 88.64 88.82 90.45 89.02 84.87 88.32 89.72 Cohen’s 39.73 45.03 46.36 52.35 45.54 38.84 45.47 50.70 MCC 41.84 46.52 48.08 53.44 46.58 44.30 47.77 52.60 REM Sleep REM Sleep REM Sleep Balanced Accuracy 75.29 81.55 82.31 86.60 77.83 77.66 82.99 88.67 70.09 85.57 84.13 89.26 Precision 90.95 93.38 92.70 94.69 91.67 93.31 92.63 94.97 90.33 93.47 92.52 94.87 Recall 90.33 93.24 91.39 94.31 90.93 93.76 90.70 94.26 91.04 91.80 89.04 93.81 F1 90.61 93.31 91.91 94.46 91.26 93.46 91.42 94.52 90.63 92.40 90.25 94.18 Cohen’s 47.59 61.83 56.59 69.06 51.59 60.63 55.18 70.06 43.76 60.09 51.95 68.78 MCC 47.72 61.85 57.27 69.23 51.80 61.06 56.35 70.54 44.03 61.23 54.38 69.61
-Appendix G: Additional Model Performance Plots
+S.3 Model Architecture
+Below is the printed output of a -class RNN model (for ; for binary models, ) with exact hyperparameter values.
+The selective SSM parameters are contained within x_proj and dt_proj.
+Softmax probabilities found using the output of this RNN model are used to represent the hypnodensities of the regular RNN.
+AnneNet(
+(net): MambaNet(
+(mamba_forward): Mamba(
+(layers): ModuleList(
+(0-2): 3 x ResidualBlock(
+(mixer): MambaBlock(
+(in_proj): Linear(in_features=10, out_features=40, bias=True)
+(conv1d): Conv1d(20, 20, kernel_size=(4,), stride=(1,), padding=(3,), groups=20)
+(x_proj): Linear(in_features=20, out_features=33, bias=False)
+(dt_proj): Linear(in_features=1, out_features=20, bias=True)
+(out_proj): Linear(in_features=20, out_features=10, bias=True)
+(norm): RMSNorm()
+(mamba_backward): Mamba(...same parameters as mamba_forward...)
+(pre_mlp): Sequential(
+(0): BatchNorm1d(140, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+(1): Linear(in_features=140, out_features=100, bias=True)
+(2): LeakyReLU(negative_slope=0.01)
+(3): BatchNorm1d(100, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+(4): Linear(in_features=100, out_features=100, bias=True)
+(5): LeakyReLU(negative_slope=0.01)
+(6): BatchNorm1d(100, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+(7): Linear(in_features=100, out_features=100, bias=True)
+(8): LeakyReLU(negative_slope=0.01)
+(9): BatchNorm1d(100, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+(10): Linear(in_features=100, out_features=10, bias=True)
+(post_mlp): Sequential(
+(0): BatchNorm1d(20, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+(1): Linear(in_features=20, out_features=100, bias=True)
+(2): LeakyReLU(negative_slope=0.01)
+(3): BatchNorm1d(100, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+(4): Linear(in_features=100, out_features=100, bias=True)
+(5): LeakyReLU(negative_slope=0.01)
+(6): BatchNorm1d(100, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+(7): Linear(in_features=100, out_features=100, bias=True)
+(8): LeakyReLU(negative_slope=0.01)
+(9): BatchNorm1d(100, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+(10): Linear(in_features=100, out_features=n, bias=True)
+The suite of decision tree architectures below are grouped by base model.
+Each configuration (separated by semicolons) is a separate model.
+All omitted parameters are set to the default values in Scikit-learn or XGBoost.
+Histogram Gradient Boosting: default parameters; max depth = 5; learning rate = 0.05; learning rate = 0.0333; learning rate = 0.01
+Random Forest: ( = 3, max depth = 7); ( = 5, max depth = 5); ( = 10, max depth = 3); ( = 10, max depth = 3)
+Extra Trees: ( = 10, max depth = 3)
+AdaBoost: = 3; = 5
+XGBoost: ( = 5, = 0.5); ( = 10, = 0.5)
+The final 3-class, 4-class, and 5-class RNN ensemble models all chose a histogram gradient boosting classifier with learning rate .
+In general, gradient boosting classifiers use many decision trees where tree is trained on the erroneous predictions of , which gradually refines performance as increases.
+For multi-class classification, regression trees compute logits for each class before applying softmax:
+As is miniscule, the RNN ensemble’s logits appear largely dominated by the initial prediction from the baseline decision tree .
+The first tree makes simple and underfitted predictions, resulting in stable but highly-uncertain probabilities such as those in Figure 3.
+S.4 Additional Model Visualizations
+UMAP projections on Figure S4 are generated using the following hyperparameters:
+metric, output_metric: euclidean
+n_neighbors: 15
+n_components: 2
+learning_rate: 1.0
+min_dist: 0.1
+spread: 1.0
+set_op_mix_ratio: 1.0
+local_connectivity: 1
+negative_sample_rate: 5
+target_n_neighbors: -1
+target_weight: 0.5
+dens_lambda: 2.0
+dens_frac: 0.3
+dens_var_shift: 0.1
+S.5 Class-Wise Evaluation Metrics
+5-Class Sleep Staging 4-Class Sleep Staging 3-Class Sleep Staging Metric (%) Regular Ensemble Regular Ensemble Regular Ensemble Wake Wake Wake Balanced Accuracy 86.47 86.19 87.93 87.87 87.97 88.27 Precision 87.64 88.44 88.34 88.67 88.24 88.82 Recall 87.64 88.04 88.32 88.70 88.12 88.85 F1 87.54 87.81 88.33 88.66 88.16 88.83 Cohen’s 73.93 74.41 75.74 76.32 75.45 72.89 MCC 74.11 75.17 75.74 76.37 75.49 76.74 N1 Sleep N1/N2 Sleep Non-REM Sleep Balanced Accuracy 67.99 69.62 71.58 73.47 81.03 85.42 Precision 86.74 87.09 75.37 75.61 81.90 85.50 Recall 81.74 79.31 75.06 75.81 80.99 85.41 F1 83.77 82.25 74.00 75.36 80.87 85.41 Cohen’s 27.21 26.46 45.39 48.25 62.01 70.83 MCC 28.61 29.06 47.03 48.74 62.90 70.92 N2 Sleep Balanced Accuracy 65.88 67.72 Precision 76.25 76.33 Recall 77.22 77.50 F1 74.80 75.82 Cohen’s 36.71 39.58 MCC 39.68 41.29 N3 Sleep N3 Sleep Balanced Accuracy 79.40 81.16 80.08 82.18 Precision 90.86 91.71 90.92 91.72 Recall 87.60 89.66 86.85 88.59 F1 88.82 90.45 88.32 89.72 Cohen’s 46.36 52.35 45.47 50.70 MCC 48.08 53.44 47.77 52.60 REM Sleep REM Sleep REM Sleep Balanced Accuracy 82.31 86.60 82.99 88.67 84.13 89.26 Precision 92.70 94.69 92.63 94.97 92.52 94.87 Recall 91.39 94.31 90.70 94.26 89.04 93.81 F1 91.91 94.46 91.42 94.52 90.25 94.18 Cohen’s 56.59 69.06 55.18 70.06 51.95 68.78 MCC 57.27 69.23 56.35 70.54 54.38 69.61
+S.6 CRNN Model Architecture
+The CRNN model reorganizes the input data into 30-second windows mirroring the 30-second windows used in visual annotation of PSG data according to AASM criteria.
+As shown in Figure S5, the CRNN model comprises:
+1)
+A convolutional component, consisting of four 1D residual blocks (see S5, right) which continually extract features and downsize the window size axis, flattening it into an embedding vector per each 30-second window.
+2)
+A recurrent component, consisting of a bidirectional (same as RNN model, S.1) three-layer Mamba block that processes the embedding vectors sequentially, followed by an element-wise multi-layer perceptron (MLP) that diminishes the feature-size axis into logits for each class.
+Below is the printed output of the layers of an -class CRNN model (for ; for binary models, ) with exact hyperparameter values.
+CRNN(
+(conv_blocks): ModuleList(
+(0): ResidualBlock(
+(net1): Sequential(
+(0): BatchNorm1d(42, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+(1): Conv1d(42, 42, kernel_size=(7,), stride=(1,), padding=(3,))
+(2): LeakyReLU(negative_slope=0.01)
+(net2): Sequential(
+(0): BatchNorm1d(42, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+(1): Conv1d(42, 64, kernel_size=(7,), stride=(2,), padding=(3,))
+(2): LeakyReLU(negative_slope=0.01)
+(mlp): Sequential(
+(0): Linear(in_features=106, out_features=64, bias=True)
+(1): LeakyReLU(negative_slope=0.01)
+(2): Linear(in_features=64, out_features=64, bias=True)
+(3): LeakyReLU(negative_slope=0.01)
+(4): Linear(in_features=64, out_features=64, bias=True)
+(5): LeakyReLU(negative_slope=0.01)
+(pool): MaxPool1d(kernel_size=2, stride=2, padding=0, dilation=1, ceil_mode=False)
+(1): ResidualBlock(
+(net1): Sequential(
+(0): BatchNorm1d(64, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+(1): Conv1d(64, 64, kernel_size=(7,), stride=(1,), padding=(3,))
+(2): LeakyReLU(negative_slope=0.01)
+(net2): Sequential(
+(0): BatchNorm1d(64, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+(1): Conv1d(64, 84, kernel_size=(7,), stride=(2,), padding=(3,))
+(2): LeakyReLU(negative_slope=0.01)
+(mlp): Sequential(
+(0): Linear(in_features=148, out_features=84, bias=True)
+(1): LeakyReLU(negative_slope=0.01)
+(2): Linear(in_features=84, out_features=84, bias=True)
+(3): LeakyReLU(negative_slope=0.01)
+(4): Linear(in_features=84, out_features=84, bias=True)
+(5): LeakyReLU(negative_slope=0.01)
+(pool): MaxPool1d(kernel_size=2, stride=2, padding=0, dilation=1, ceil_mode=False)
+(2): ResidualBlock(
+(net1): Sequential(
+(0): BatchNorm1d(84, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+(1): Conv1d(84, 84, kernel_size=(7,), stride=(1,), padding=(3,))
+(2): LeakyReLU(negative_slope=0.01)
+(net2): Sequential(
+(0): BatchNorm1d(84, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+(1): Conv1d(84, 104, kernel_size=(7,), stride=(2,), padding=(3,))
+(2): LeakyReLU(negative_slope=0.01)
+(mlp): Sequential(
+(0): Linear(in_features=188, out_features=104, bias=True)
+(1): LeakyReLU(negative_slope=0.01)
+(2): Linear(in_features=104, out_features=104, bias=True)
+(3): LeakyReLU(negative_slope=0.01)
+(4): Linear(in_features=104, out_features=104, bias=True)
+(5): LeakyReLU(negative_slope=0.01)
+(pool): MaxPool1d(kernel_size=2, stride=2, padding=0, dilation=1, ceil_mode=False)
+(3): ResidualBlock(
+(net1): Sequential(
+(0): BatchNorm1d(104, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+(1): Conv1d(104, 104, kernel_size=(7,), stride=(1,), padding=(3,))
+(2): LeakyReLU(negative_slope=0.01)
+(net2): Sequential(
+(0): BatchNorm1d(104, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+(1): Conv1d(104, 126, kernel_size=(7,), stride=(2,), padding=(3,))
+(2): LeakyReLU(negative_slope=0.01)
+(mlp): Sequential(
+(0): Linear(in_features=230, out_features=126, bias=True)
+(1): LeakyReLU(negative_slope=0.01)
+(2): Linear(in_features=126, out_features=126, bias=True)
+(3): LeakyReLU(negative_slope=0.01)
+(4): Linear(in_features=126, out_features=126, bias=True)
+(5): LeakyReLU(negative_slope=0.01)
+(pool): MaxPool1d(kernel_size=2, stride=2, padding=0, dilation=1, ceil_mode=False)
+(rnn): MambaNet(
+(mamba_forward): Mamba(
+(layers): ModuleList(
+(0-2): 3 x ResidualBlock(
+(mixer): MambaBlock(
+(in_proj): Linear(in_features=126, out_features=504, bias=True)
+(conv1d): Conv1d(252, 252, kernel_size=(4,), stride=(1,), padding=(3,), groups=252)
+(x_proj): Linear(in_features=252, out_features=40, bias=False)
+(dt_proj): Linear(in_features=8, out_features=252, bias=True)
+(out_proj): Linear(in_features=252, out_features=126, bias=True)
+(norm): RMSNorm()
+(mamba_backward): Mamba(...same parameters as mamba_forward...)
+(mlp): Sequential(
+(0): BatchNorm1d(252, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+(1): Linear(in_features=252, out_features=100, bias=True)
+(2): LeakyReLU(negative_slope=0.01)
+(3): BatchNorm1d(100, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+(4): Linear(in_features=100, out_features=100, bias=True)
+(5): LeakyReLU(negative_slope=0.01)
+(6): BatchNorm1d(100, eps=1e-05, momentum=0.1, affine=True, track_running_stats=True)
+(7): Linear(in_features=100, out_features=100, bias=True)
+(rnn_activation): LeakyReLU(negative_slope=0.01)
+(mlp): Sequential(
+(0): Linear(in_features=226, out_features=64, bias=True)
+(1): LeakyReLU(negative_slope=0.01)
+(2): Linear(in_features=64, out_features=64, bias=True)
+(3): LeakyReLU(negative_slope=0.01)
+(4): Linear(in_features=64, out_features=16, bias=True)
+(5): LeakyReLU(negative_slope=0.01)
+(6): Linear(in_features=16, out_features=n, bias=True)
+S.7 Additional Model Performance Plots
+![[Uncaptioned image]](figures/unhealthy-subset.png)
 Metric (%) Classes Balanced Accuracy Precision Recall F1 Cohen’s MCC 5 63.56 70.91 63.34 65.30 51.43 52.60 4 73.68 75.40 73.15 73.73 60.11 60.46 3 83.13 84.02 83.13 83.38 71.28 71.41
+S.8 Accelerometry Signals
+We aggregate our pre-processed accelerometry signals from the ANNE One chest module into Eucliean Norm Minus One (ENMO)1 and z-angle2.
+On Figures S12, S13, S14, S15, and S16, we select five representative full night recordings of healthy individuals from our study cohort and plot them against the ground truth hypnogram to show how accelerometry informs sleep stage transitions at the individual level.
+We also show power spectral density (PSD) heatmaps stratified by sleep stage for each recording.
+S.9 Accelerometry Signal Ablation Study
+Metric (%) Classes Feature Set Balanced Accuracy Precision Recall F1 Cohen’s MCC Accel.
++ Temp.
+63.85 70.68 63.35 65.03 51.97 53.12 Accel.
+64.60 71.15 64.18 65.64 52.97 54.12 5 Accel.
+(0.5 Hz High-Pass) 57.40 67.00 60.84 62.31 48.21 49.12 [1pt/1pt] Accel.
++ Temp.
+73.03 74.46 71.52 72.10 58.70 59.30 Accel.
+73.99 75.27 72.64 73.14 60.21 60.77 4 Accel.
+(0.5 Hz High-Pass) 67.23 72.00 68.98 69.55 54.98 55.65 [1pt/1pt] Accel.
++ Temp.
+82.78 84.26 83.18 83.51 71.71 71.56 Accel.
+83.08 84.82 83.76 84.08 72.47 72.61 3 Accel.
+(0.5 Hz High-Pass) 76.11 81.48 80.13 80.60 66.45 66.64
+S.10 Feature Importance by Sensor